Precision Gas Sampling (PGS) Instrument Log

 

PGS    C1       CF Lamont, OK 60M Tower

 

Date

Technician

Comments

10/2/15

MG

Could not open the camera on the web page. Power cycled camera and reloaded the web page all is well

 

4/9/2015

PD

PGS shed at the base of the 60M tower is being replaced with a 20' ISO container.  Crane is scheduled to remove the current PGS shed today.  Crane lifted PGS shed onto trailer and secured.  RTS at 1610 GMT.

3/30/15

PD/KT

Radon Gas Analyzer (RGA) instrument is to be removed from service.  Shut system down at 1520 GMT.  Removed computer, calibrator and all components from end of instrument enclosure.  No model or SN are found on calibrator or main instrument enclosure. Laptop computer is a Dell Latitude D505, Service Tag 90CFH71, Express Service Code 19611918541, CN-OH2049-48643-54M-5338.

2/7/13

PAD

CO2flux computer (aka PGS computer) that collects data from 4M met system failed.  Mentor sent replacement computer and computer dept was able to recover and copy all relevant files to replacement computer.  Mentor requests computer installed.  Assisted Windows administrator installing computer at 1930 GMT.  Administrator loaded necessary drivers for Keyspan serial-usb convertors and Picarro shed Nikon camera.  Computer will now communicate with camera and 4M met datalogger.  Some settings on PC208 for datalogger were not known.  Notified Mentor computer was installed and asked that he check datalogger settings.  Dell Optiplex 745, Model DCCY, SN F14KRD1, LBLDOE 6658330.  RTS at 2100 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  Power was restored at 2115 GMT after electrical upgrades were complete. 

11/14/12

KT

Valve on RGA instrument was left closed and fan shut down after power was restored on 11/13 to allow particles to drop from measurement chamber.   Valve was opened and fan was powered back on today.  RTS at 2055 GMT.

11/13/12

PAD/KT

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  Valve on RGA instrument was left closed and fan shut down after power was restored to allow particles to drop from measurement chamber.  Unit will be RTS tomorrow.  RTS at 2130 GMT.

4/19/12

PAD

Changed desiccant at 1458 GMT.

3/12/12

PAD

RGA inlet pipe that was separated approximately 40M above the ground was repaired by contractor on 3/8/12.  Mentor has provided plans to alter horizontal inlet pipe from bottom of tower to RGA inlet and requests this work performed since inlet pipe has been repaired on tower.  Assisted facilities department in altering horizontal pipe as instructed by mentor.  Pipe was raised off ground to allow larger water trap below horizontal pipe.  Pipe was sloped down towards tower in case any water did enter horizontal pipe.  Pictures were taken before/after alteration and mentor was notified when upgrade was completed.  RTS at 2002 GMT.

3/8/12

PAD

RGA inlet pipe has been separated approximately 40M above ground for some time.  Contractor is on site today to climb tower and fix broken pipe.  Shut off power to tower at base at 2023 GMT.  Contractor climbed tower and repaired broken collar on RGA inlet tubing.  Contractor inspected  and took pictures of inlet tubing connectors from snorkel located on top of tower to ground.  No other problems were found.  Power was restored to tower at 2257 GMT.  RGA was not shut down during this time but is now sampling air from 60M level.  RTS at 2257 GMT.

3/8/12

PAD

At 2000 GMT, observed RGA flow had dropped to near zero around 1300 GMT.  RGA inlet pipe has been separated approximately 40M above ground for some time.  At 2000 GMT, disconnected horizontal pipe from base of tower and barrels at other end.  Laid pipe on it's side and drained small amount of water out but not enough to block flow.  Removed cap from end of pipe and dead bird came out of pipe.  Bird must have entered inlet pipe at separation.  Reconnected pipe.  Contractor on site today to fix broken pipe.  See additional CM report for details.  RTS at 2015 GMT.

1/19/12

PAD

Flow had dropped to zero for RGA.  Mentor was on site on 11/14/12 and found water had filled horizontal pipe between tower and instrument.  Tubing connecting pipe to barrels was broken at that time.  Removed water from pipe on 1/18/12 but did not have replacement tubing at that time.  Installed tubing at 1958 GMT.  Pipe is still broken on the tower.  Mentor is aware of this and pipe is scheduled to be fixed in February 2012.  RTS at 2013 GMT.

7/26/11

PAD

*PGS system is being shut down and removed.  System is being replaced with Picarro system.  PGS system was shut down on 7/21/11.  At 1400 GMT, removed PGS rack, control boxes, and pumps.  Installed Picarro system using pumps and refrigerator from PGS system.  Took all removed items from PGS to shipping to send to LBNL.  RTS at 2200 GMT.

7/21/11

PAD

PGS system is being shut down and removed from service.  It will be replaced with the Picarro system.  Powered down instrument rack at 1628 GMT.  Left refrigerator and vacuum pumps running as per mentor's request.  Mentor will be on site later this month to remove PGS system.  RTS at 1643 GMT.

7/11/11

PAD

Notified by mentor he is unable to communicate with PGS computer.  Computer power indicator light is flashing.  Powercycled computer and computer started correctly.  Able to connect with PGS datalogger and data is now being transferred.  Mentor is now able to remotely access computer.  There were other unrelated instruments that failed at the same time.  Suspect power outage or surge.  RTS at 1845 GMT.

6/28/11

PAD

Water in line between adjustable flow valve and back pressure indicator gauge.  Removed line at gauge and allowed water to blow out.  Reconnected line.  RTS at 1631 GMT.

6/14/11

PAD

PGS vacuum pot pump has failed.  R/R vacuum pump with a spare pump for CO system.  System is now operating at 24 inHg.  RTS at 1828 GMT.

4/19/11

PAD

#2, 3, and 4 discharge tubes coming out of refrigerator are plugged.  At 1800 GMT, R/R copper drain tubes for #1, 2, 3, and 4.  All drain tubes will now flow air and allow water to be expelled from traps.  RTS at 1825 GMT.

3/29/11

PAD

#2 sample line discharge tube is blocked.  Cleared with canned air at 1918 GMT.

3/24/11

PAD

RGA was shut down on 3/23/11 for background counts and needs to be restarted.  Opened inlet line valve and powered up instrument at 1432 GMT.

3/23/11

PAD

Mentor requests RGA unit shut down for background counts.  Shut down instrument and closed inlet line valve at 1400 GMT.

3/22/11

PAD

Signal wire between the calibration source and the controller located in the Radon Gas Analyzer (RGA) controller had been cut by rodent.  Removed signal wire and took to lab.  Installed repaired wire at 1600 GMT.  RTS at 1620 GMT.

3/10/11

PAD

Notified by mentor PGS IRGA pressure was lower than normal and dewpoint has risen to -6.3.  Desiccant had been changed on 3/1/11.  Checked connections and cap.  Connections were tight and got very little out of tightening cap.  Pressure increased to 792.67 and dewpoint is slowly starting to drop.  RTS at 2012 GMT.

3/9/11

PAD

Request by mentor to replace NOAA sampler.  Old Sherpa Model 70, SN 1-004. New SN 1-001.  RTS at 1952 GMT.

3/9/11

PAD

Reference cylinder shows 250 psi and needs replaced.  R/R cylinder at 1916 GMT.  Checked for leaks and purged line.  Old XCO14813B, New CC94145.  New cylinder shows 1940 psi.  RTS at 1936 GMT.

2/23/11

PAD

Pressr_4 = 7.79 l/min on 60M inlet.  Vacuum pump is failing.  No replacement pump of the same model on site.  Shut down and removed vacuum pot pump from PGS system that is the same model.  Installed temporary pump on vacuum pot.  Installed removed pump on 60M inlet.  Flow is now normal at 17.8 l/min.  Old- cannot access model and SN.  New- model UN010 ANI, SN 1/1026011.  RTS at 1447 GMT.

2/10/11

PAD

IRGA temperature is 37.1C.  Temperature controller shows:  Bad input.  Powercycled Power Supply Box.  IRGA temperature is slowly increasing.  RTS at 1949 GMT.

2/8/11

PAD

#3 sample line discharge tubing is blocked.  Cleared with canned air at 1643 GMT.

2/3/11

PAD

PGS IRGA enclosure temperature has been sporadic and has now dropped to 38.7C.  Temperature controller on Power Supply Box shows:  Bad Input.  Powercycled Power Supply Box.  Bad input message is gone and enclosure temperature is slowly warming.  RTS at 2041 GMT.

12/1/10

PAD

AC power to CF will be shut down today for ongoing construction.  Instruments affected are PGS, PICARRO, CO, Radon Gas Analyzer, and CO2 Flux.  Restarted systems at 1909 GMT.

11/22/10

PAD

AC power will be interrupted for construction near the main compound.  Need to gracefully shut the instrument down prior to power outage.  Instruments affected are PGS, CO, RGA, and CO2flux.  Powered instrument down at 1420 GMT.  Restarted instrument after AC power was restored.  RTS at 1800 GMT.

11/18/10

PAD

AC power needed to be secured to drill pier holes for new construction near the RCF trailer.  Need to gracefully shut the instrument down prior to power outage.  Powered system down at 1420 GMT. Shut down PGS, Radon Gas, CO2 flux and CO computers. Shut down PGS IRGA rack and all pumps located in PGS enclosure. Powered down Radon Gas Analyzer unit. Restarted systems after AC power was restored. RTS at 1750 GMT.

11/15/10

PAD

PGS reference cylinder is down to 300 psi and needs to be replaced.    Shut off cylinder and R/R.  Checked for leaks at cylinder connector and purged line.  Old XC006452B, 366.1 ppm.  New XC014813B, 370.2 ppm.  RTS at 1825 GMT.

11/10/10

PAD

AC power will be interrupted for construction near the main compound.  Need to gracefully shut the instrument down prior to power outage.   Powered system down at 2118 GMT.  Shut down PGS, Radon Gas, CO2 flux and CO computers.  Shut down PGS IRGA rack and all pumps located in PGS enclosure.  Powered down Radon Gas Analyzer unit.  Restarted systems after AC power was restored.  RTS at 2336 GMT.

9/7/10

PAD

Notified by mentor 60m pump, #4, had failed on 9/6/10 at 1400 GMT.  R/R *pump type UN010 ANI.  RTS at 2130 GMT. 

*Old - 1/1041919   *New - 1/1195187

8/6/10

JEM

Radon Gas Analyzer was shut down on 8/4/10 to allow farmer to burn surrounding field.  Instrument was left powered down to allow particles to fall off chamber and needs to be started.  Powered up instrument and opened inlet valve at 1319 GMT.

8/4/10

PAD

Farmer intends to burn field PGS trailer is located in.  System needs to be shut down to avoid smoke being drawn in.  Stopped programs and shut down vacuum pumps at 0034 GMT.  Systems shut down were PGS, CO, and Radon Gas Analyzer.  All systems except Radon Gas Analyzer were restarted at 0229 GMT.

7/23/10

PAD

Reference cylinder is down to 290 psi and needs to be replaced.  R/R cylinder at 1551 GMT.  Checked for leaks at cylinder and purged line.  Old CC157178/ 366.4 ppm CO2.  New XC006452B/ 366.1 ppm CO2.  RTS at 1611 GMT.

6/22/10

PAD

#2 drain tube in PGS system is blocked.  Cleared blockage with canned air at 1446 GMT. 

6/11/10

PAD

Changed PGS desiccant at 2020 GMT.

5/27/10

PAD

Mentor request CO pump R/R with rebuilt pump provided by mentor.  Shut off pump and R/R at 1843 GMT.  RTS at 1858 GMT.  Old SN 1/1041267.  New SN 1/1698151.

6/1/10

PAD

#2 and #3 sample line discharge tubes are obstructed.  Water is visible in #2 and #3 flowmeters.  Removed tubes and cleared obstruction with canned air.  Removed inlet and outlet lines on flowmeters and removed water with canned air.  RTS at 1625 GMT.

5/27/10

PAD

At 1522 GMT, observed #4 sample line inlet pressure at 1.78 psi.  Graphs show pressure dropped from 12 psi to current level on 5/26/10 around 1000 GMT.  Troubleshooting found #4 pump is running but unable to maintain adequate pressure to sampling system.  At 1824 GMT, R/R pump rack assembly.   #1-4 sample line pressures are all in the 17-18 psi range.  RTS at 1839 GMT.  Old SN 1/1195183, 1/1198184, 1/1195185, 1/1195186.  New SN 1/952959, 1/1041922, 1/1041921, 1/1041919.

5/17/10

PAD

CF lost power on 5/16/10 between 0952-1504 GMT.  At 1557 GMT, rebooted RGA computer to establish network connection.  RTS at 1612 GMT.

5/17/10

PAD

CF lost power on 5/16/10 between 0952-1504 GMT.  Logged onto PGS  computer at 1233, connected with PGS and CO2Met dataloggers to transfer data.  Reset Keyspan serial to USB converters.  Reset camera.  RTS at 1253 GMT.

4/29/10

PAD

Changed desiccant at 2002 GMT.

3/30/10

PAD

PGS reference cylinder is at 290 psi and needs to be replaced.  Shut off flow and R/R cylinder at 1541.  Checked for leaks and purged line.  New cylinder pressure is 2020 psi.  Old  CC149050, 370.3 ppm.  New CC157178, 366.4 ppm.  RTS at 1610 GMT.

2/10/10

PAD

Calibrator was removed on 2/9/10 to troubleshoot flow issue and needs to be installed.  Installed calibrator at 1846 GMT and started calibration manually.  Set flow at 50 cc/min.  RTS at 1914 GMT.

2/9/10

PAD

Notified by mentor instrument is not getting flow from calibrator during calibration cycle.  Removed calibrator at 2021 GMT and took to lab.  Troubleshooting indicates adjustable flowmeter downstream of 12V pump was closed likely due to vibration.

1/21/10

PAD

Span gas 1 cylinder is down to 300 psi and needs to be replaced.  R/R cylinder at 1630 GMT.  New cylinder shows 1860 psi.  Old SG9199683.  New CC157152.

12/8/09

PAD

Mentor requests PGS pumps swapped with spare pump rack supplied.  Shut down pumps, R/R with replacement pumps for 2,4,25, and 60m.  Old SN 1/1041919, 1/952959, 1/1041921, 1/1041922.  New SN 1/1195183, 1/1195184, 1/1195185, and 1/1195186.  RTS at 2132 GMT.

12/3/09

PAD

PGS reference cylinder is down to 260 psi. and needs replaced. R/R reference cylinder, checked for leaks, and purged line.  Old CC176843/ 370.7 ppm.  New CC149050/ 370.3 ppm./ 1880 psi. cold.  RTS at 2210 GMT.

11/10/09

PAD

#4 sample line discharge tubing is blocked.  Removed tubing and cleared with canned air at 1624 GMT. 

11/3/09

PAD

PGS computer is unable to access network.  CO and RGA computers will access.  Shut down/ restarted PGS computer.  It will now access network.  RTS at 1715 GMT.

8/18/09

PAD

#4 sample line drain tube was blocked.  Removed drain line and cleared with canned air.  Removed inlet and outlet lines at flowmeter and cleared water with canned air.  RTS at 2105 GMT.

8/6/09

PAD

Reference cylinder is at 300 psi. and needs replaced.  R/R reference cylinder, checked for leaks and purged line.  Old CC175907/ 368.8 ppm.  New CC176843/ 370.7 ppm.  New cylinder pressure = 2070 psi.  RTS at 2140 GMT.

6/16/09

PAD

#3 sample line drain tube is plugged.  There is water in #3 flowmeter in sample rack.  R/R #3 drain tube and disconnected inlet and outlet lines from flowmeter.  Cleared water with canned air.  RTS at 1605 GMT.

6/10/09

PAD

Radon Gas Analyzer was shut down on 6/9/09 for annual measurement of accumulated background counts and needs to be restarted.  Powered up instrument and opened inlet line valve at 1824 GMT.

6/9/09

PAD

Mentor requests Radon Gas Analyzer shut down for annual measurement of accumulated background counts.  Shut down instrument and closed inlet line valve at 1829 GMT.  Instrument will remain shut down for 24 hours. 

5/29/09

PAD

PGS vacuum tank pump is still holding -20 inHg, but is starting to fail and needs replaced.  At 1437 GMT, shut down and removed Thomas pump, model 2107CA18, and installed KNF pump, model UN010 ANI.  Old SN 000011.  New SN 1/1026011.

5/29/09

PAD

Mentor requests rack containing PGS pumps R/R with rebuilt pumps.  Pump #2 in replacement rack is not working and needs to be replaced with spare pump.  At 1347 GMT, shut down and removed pump #4.  R/R pump # 2 in replacement PGS rack with removed #4 pump.   New SN 1/952959.  Shut down pumps 1-3 at 1437 and R/R pump rack assembly.  Restarted pumps 1-4 and checked flows.  RTS at 1452 GMT.

5/28/09

PAD

Mentor requests rack containing PGS pumps R/R with rebuilt pumps.  Shut pumps down at 2005 GMT and installed new pumps.  Pumps #1 and 2 would not start up.  Installed original rack for PGS until pump problems could be corrected.  Troubleshooting found #1 pump on/off switch only has 55V going to pump.  Bypassed individual switch so main on/off switch is now controlling pump #1.  Pump #2 appears to have been damaged in shipping.  Will R/R pump #2 with spare tomorrow morning and install new rack.

5/5/09

PAD

#3 sample line discharge tube is blocked.  Cleared blockage with canned air.  RTS at 1530 GMT.

4/17/09

PAD

Reference cylinder is down to 280 psi and needs to be replaced. R/R reference cylinder, checked for leaks, and purged line.  Old CC140047, New CC175907/ 368.8 ppm/ 1950 psi.  RTS at 2035 GMT.

4/16/09

PAD

#3 sample line discharge tube is blocked.  Cleared block with canned air at 2002 GMT. 

4/16/09

PAD

R/R PGS desiccant canister at 2012 GMT.  RTS at 2027 GMT.

2/18/09

PAD

Notified by mentor that #4 sample flow is low.  Pump is operating and no leaks were found.  Installed spare pump and now Pressr_4 = 13.1.   Removed SN 1/1026010 and installed SN 1/952959.  RTS at 1905 GMT.

1/8/09

PAD

Notified by mentor values for reference gas on PGS IRGA system have been noisy since 12/30/08.  Mentor requests computer rebooted, IRGA and datalogger power cycled.  Rebooted computer and powercycled IRGA rack at 2059 GMT.  RTS at 2114 GMT.

12/30/08

PAD

Reference cylinder needs to be replaced.  Removed cylinder CC16121, 370.0 ppm CO2, and installed cylinder CC140047, 370.0 ppm.  Old cylinder shows 240 psi and new cylinder shows 1850 psi.  Purged line and checked for leaks.  RTS at 1649 GMT.

11/25/08

PAD

#1 sample drain line is blocked.  Cleared obstruction with canned air.  RTS at 1706 GMT.

10/31/08

PAD

Notified by mentor 25M level sample is spiking in CO2 concentration values.  Mentor suspects leak in pump or inlet line causing system to sample room air.  T/S found inlet tubing had broken at compression fitting on pump #3.  Fixed tubing and reconnected.  RTS at 1626 GMT.

10/30/08

PAD

Notified by mentor 25M level sample is spiking in CO2 concentration values.  Mentor suspects leak in pump or inlet line causing system to sample room air.  At 1405 GMT, monitored system and observed #3 sample reading approximately 100 ppm CO2 higher concentration than other 3 levels.  T/S including switching lines at 1417 GMT with #2 and #3 pump.  Higher CO2 readings remained with #3 pump.  Switched lines back at 1427 GMT.  At 1925 GMT, R/R pump #3.  Monitored concentrations and #3 sample is still 50 ppm CO2 higher than other levels.  Switched #3 and #4 sample lines at 2108 GMT.  Higher CO2  levels remained with #3 pump.  Switched lines back at 2138 GMT.  Suspect faulty quick connect on inlet connector of pump #3.  Troubleshooting at this time.  RTS at 2150 GMT.

10/20/08

PAD

Mentor requests PGS pump rack assembly swapped with replacement pumps provided.  R/R pump rack assembly for PGS system at 1414 GMT.  RTS at 1429 GMT.

10/20/08

PAD

PGS computer has power but nothing is shown on monitor.  Checked connections and verified CO2 flux computer was working.  Power cycled PGS computer and it started correctly.  RTS at 1418 GMT.

8/26/08

PAD

Changed IRGA desiccant at 1610 GMT.  RTS at 1625 GMT. 

8/14/08

PAD

PGS standard cylinder, 460 ppm, is dropping 10 psi. between weekly PM checks.  Found leak at cylinder/ regulator connection.  Removed regulator from cylinder and retaped with teflon.  Reconnected regulator and checked for leaks.  RTS at 1935 GMT.

7/30/08

PAD

Mentor requests PGS calibration cylinders replaced with cylinders provided.  Shut off valves to calibration cylinders at 1840 GMT.  Removed cylinders CA05655, CA05687, CA05679, and CA05658.  Installed cylinders CA07525, CA07537, CA07414, and CA07514.  These cylinders have CO2 concentrations of 346.45/ 377.42/ 406.95/ and 464.06 ppm respectively.  Checked for leaks at regulator/ cylinder connection.  Opened outlet flow valve at 2001 GMT and set low pressure side of regulators to 15 psi.  Took empty cylinders to shipping/ receiving.

7/13/08

PAD

CF lost power on 7/13/08 between 0028 and 0137 GMT. 

7/9/08

PAD

CF experienced power loss from 0745 to 1358 GMT.  Logged onto PGS computer.  Reset serial/USB converter to connect with datalogger.  Powered on CO2Flux computer and started Sonic1, 2, and 3 programs.  Started CO computer.  Rebooted Radon Gas Analyzer computer to reestablish comms.  Reset camera date and time.  RTS at 1515 GMT.

6/24/08

PAD

#3 sample line discharge tube is blocked.  Small amount of water is in #3 flowmeter.  At 1556 GMT, removed discharge tubing and cut off last 2" to try to remove suspected trouble spot .  Removed inlet and outlet lines from #3 flowmeter and dried flowmeter with canned air.  Connected flowmeter lines and discharge tubing and verified system is operational.  RTS at 1611 GMT.

6/17/08

PAD

Notified by mentor data files are not usable.  Mentor requests computer rebooted.  Suspect power outages on 6/16/08 may have caused corruption.  Rebooted computer at 1915 GMT.  RTS at 1930 GMT.

6/17/08

PAD

Notified by mentor he cannot remotely access Radon Gas Analyzer computer.  Suspect network connection problem between computer and network hub.  At 1532 GMT, tried to repair network connection using computer software but was not successful.  Stopped program and rebooted computer.  Network connection is now established.  RTS at 1547 GMT.

6/9/08

PAD

CF experienced extended power outage early on 6/9/08.  PGS shed has lost comms.  Reset media convertors located in PGS shed and AOS trailer.  RTS at 1524 GMT.

6/10/08

PAD

Notified by mentor of fluctuation in PGS IRGA flow values.  Troubleshooting traced problem to loose connection or short in data cable between IRGA enclosure bulk head and datalogger enclosure bulkhead.    Shut down PGS system and removed cable.  Took cable to lab where Mark Klassen found pins in the female connector of the cable were not seated correctly.  After repairs, I installed cable and restarted PGS system at 1830 GMT.  RTS at 1830 GMT.

6/2/08

PAD

CF lost power on 6/1/08.  Communications were lost between PGS shelter and AOS trailer.  Reset media converters located at PGS shelter and AOS trailer.  RTS at 1320 GMT.

5/22/08

PAD

#3 sample line discharge tube is plugged.  Cleared obstruction with canned air.  RTS at 1505 GMT.

5/13/08

PAD

Reference cylinder pressure is 260 psi. and needs replaced.  At 2000 GMT, shut off valve and removed reference cylinder CC12133, 370.5 ppm., and installed cylinder CC201086, 369.8 ppm. CO2.  Checked for leaks and purged line.  New cylinder pressure shows 2,100 psi.  RTS at 2030 GMT.

5/1/08

PAD

Radon GA - Notified by mentor he cannot remotely access computer.  Found computer is running.  Stopped program and rebooted computer.  Computer will now access network.  Notified mentor.  RTS at 1458 GMT.

4/26/08

PAD

CF experienced extended power outage on 4/26/08 at 0034 GMT.  Power was restored at 0459 GMT.

4/29/08

PAD

CF experienced extended power outage on 4/26/08 at 0034 GMT.  Power was restored at 0459 GMT.  At 1534 GMT, found computer running but I had to log on.  Power cycled Keystone serial/ USB converter.  Computer will now communicate with datalogger.  RTS at 1623 GMT.

4/8/08

PAD

Power outage at CF between 0709 and 0910 GMT.  Reset media converters located in PGS and AOS trailers at 1430 GMT.

3/19/08

PAD

Mentor requests PGS pump rack assembly changed out with replacement rack.  At 2020 GMT, shut down and removed *pumps, SN 1/1026010, 1/1026011, 1/1026012, 1/1026013.  Installed rack containing pumps, SN 1/1041919, 1/1041920, 1/1041921, 1/1041922.  Started pumps and adjusted pressures.  RTS at 2045 GMT.

2/26/08

PAD

#2 flowmeter is showing 0 flow.  Suspect ball is stuck at bottom.  Removed inlet line and used canned air to dislodge ball.  Reconnected inlet line.  Flowmeter is now showing flow.  RTS at 1907 GMT.

2/21/08

PAD

Sample line #3 is currently connected to test cylinder.  Mentor request sample line removed from cylinder and connected to PGS sample rack.  At 2045 GMT, disconnected line from cylinder and connected to PGS sample rack.  RTS at 2100 GMT.

2/5/08

PAD

Span 1 cylinder shows 290 psi. and needs to be replaced.  Removed cylinder, (CC210699, 0.118 ppm CO.), and installed cylinder, (CC210306, 0.107 ppm CO).  New cylinder shows 1850 psi.  Checked for leaks and opened valve.  RTS at 2041 GMT.

1/23/08

PAD

Mentor request reference cylinder replaced.  Cylinder is showing 275 psi.  At 1703 GMT, shut off cylinder valve and removed reference cylinder, (XC011414B, 367.6 ppm CO2), and installed cylinder, (CC12133, 370.5 ppm CO2).  At 1712 GMT, opened cylinder valve, purged line, and checked connections for leaks.  New cylinder shows 1860 psi.  Notified instrument mentor.  RTS at 1726 GMT.

1/7/08

PAD

Mentor request cylinder currently connected to #1 sample line for testing connected to #3 sample line and #1 sample line connected to 2m inlet line.  At 2200 GMT, removed test line from #1 sample line and connected to #3 sample line.  Connected #1 sample line to 2m inlet line.  RTS at 2215 GMT.

11/5/07

PAD

Notified by mentor #2 sample flow is 0.  Found #2 sample outlet line from pump broke off at compression nut.  Repaired line and flow returned to normal.  Notified mentor.  RTS at 1910 GMT.

11/1/07

PAD

Computers, comm boxes, and instruments need to be moved to new UPS installed in PGS shed.  Stopped 4m, 25m, and 60m Sonic programs.  Shut down PGS and CO2 Flux computer.  Moved computer power cables and comm power supplies to new UPS unit.  Restarted computers and Sonic programs.  RTS at 1951 GMT.

10/17/07

PAD

Notified by mentor he is unable to remotely access PGS computer.  Heavy thunderstorms at CF. At 1856 GMT, reset media converter boxes in PGS and Aerosol buildings.  Comms are now restored.  Notified mentor.  RTS at 1911 GMT.

10/5/07

PAD

Found #2 sample discharge line plugged.  Removed drain line and cleared obstruction with canned air.  Removed inlet and outlet lines from flowmeter and blew out water trapped in flowmeter.   RTS at 1820 GMT.

10/3/07

PAD

Mentor request PGS pump assembly changed out with replacement pump rack.  R/R PGS pump rack assembly at 1226 GMT.  RTS at1241 GMT.

10/3/07

PAD

CF experienced loss of power on 10/2/07.  Power has been restored and instrument needs RTS.  At 1208 GMT, turned on breakers to PGS shed.  UPS that supplies power to PGS, Radon, and CO computers shorted out.  R/R UPS with XL-1000 UPS and restored power.  Reestablished comms to PGS shed.  RTS at 1338 GMT.

10/2/07

PAD

CF experienced loss of power at 1945 GMT.  Utilities company informed us outage could last for hours.  Management decided to shut instruments down to prevent problems at start up.  At 2009 GMT, found computer and instrument had already shut down.  Shut off breakers to instruments and UPS.  RTS at 2024 GMT.

8/16/07

PAD

R/R desiccant and installed quick connectors supplied by mentor to desiccant canisters at 1952 GMT.  RTS at 2007 GMT.

7/26/07

CAM

West booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1354 GMT.  Raised booms and RTS at 1900 GMT.

7/12/07

PAD

Desiccant needs changed on instrument.  At 2035 GMT, R/R desiccant cartridge. RTS at 2050 GMT.

6/27/07

PAD

Mentor request PGS reference cylinder replaced.  It is down to 220 psi.  Dave Billesbach is on site and brought reference cylinder from AirGas.  At 1516 GMT, shut off reference cylinder, XC029941C 370.2 ppm, and installed cylinder XC007637B 344.4 ppm.  Purged line from cylinder to PGS rack and checked for leaks.  RTS at 1536 GMT. 

6/19/07

PAD

#1 sample line discharge tube is blocked.  Removed tubing from collection flask and cleared with canned air.  RTS at 1950 GMT.

6/14/07

PAD

#2 sample line discharge tubing is blocked.  At 1958 GMT, removed tubing and cleared blockage with canned air.  RTS at 2009 GMT.

6/7/07

PAD

PGS system has been experiencing sporadic flow.  Mentor requests some troubleshooting to determine cause for sporadic flow and R/R mass flow controller.  At 1511 GMT, started troubleshooting flow issue with Dave Billesbach via phone.  Disconnected outlet line at mass flow meter and flow was low on #1 and #2 sample line.  Reconnected outlet line and switched #1 and #3 inlet and outlet line at pumps.  #3 line on #1 pump= 99ccm.  #1 lines on #3 pump= 89 ccm.  At 1600 GMT, shut off reference cylinder valve and shut down PGS rack.  R/R mass flow controller as per instructions.  At 1642 GMT, powered up rack and purged reference gas line.  While switching inlet and outlet lines on pumps #1 and #3, found head bolt on top of pump #1 was backed out.   Tightening bolts on #1 pump raised Pressr_1 from 13.56 psi to 14.43 psi.  Tightened all head bolts on other three pumps.  Flow on all pumps is steady at ~99ccm.  RTS at 1701 GMT.

6/5/07

PAD

Notified by mentor flow to IRGA in PGS rack is sporadic and low.  MasPl shows 77.0 ccm. This happened after power outage at CF last week.  Mentor advised power cycling PGS rack to see if flow controller would reset.  Power cycled PGS rack at 2023 GMT.   MasPl shows 99.22 ccm.  Left voice mail with instrument mentor.  Will continue to monitor.  RTS at 2038 GMT.

4/11/07

PAD

Mentor request PGS IRGA removed and sent for repairs.  At 1422 GMT, Mark Klassen and I shut off power to data logger box, valve box, isotope sampler box, and main power supply box.  Shut off reference cylinder valve.  Removed lines and connections as instructed and removed IRGA from instrument rack.  Took instrument to shipping dept.  RTS at 1502 GMT.

3/20/07

PAD

Line #4 drain tube is plugged.  Removed tube and obstruction from fitting on collection flasks.  RTS at 2103 GMT.

3/2/07

PAD

PGS system reference cylinder needs to be replaced.  Removed cylinder SG9164659BAL, 368.7 ppm CO2, and installed cylinder XC029941B, 370.2 ppm CO2.  Checked regulator connector for leaks and purged line for one minute.  Took empty cylinder to rack for pickup by AirGas.  RTS at 2122 GMT.

3/1/07

PAD

Mentor request PGS pump rack changed out with spare rack received at CF. Shut down pumps at 2046 GMT.  Disconnected lines and removed *pump rack containing pumps, (Model UN010 ANI, SN's 1/593121, 1/713617, 1/713616, 1/593124).  Installed pump rack assembly containing pumps, (Model UN010, SN's 1/713615, 1/640085, 1/593123, 1/593122).   Started pumps and adjusted pressure readings on display to 13.0.  RTS at 2101 GMT.

2/13/07

PAD

Desiccant needs to be changed for PGS system IRGA.  R/R desiccant at 1951 GMT.  RTS at 2006 GMT.

1/2/07

PAD

Vacuum pump in PGS system has been failing and is down to 8.0 in. -Hg.  Removed Dayton Speedaire pump model 4Z026, SN 0982 and plumbed and installed Thomas model 2107CA18 3, SN 000011, a spare pump found in PGS shed.  Switched out pumps at 1745 GMT.  New pump shows vacuum of 20 in. -Hg.  RTS at 1800 GMT.

12/19/06

PAD

#2 sample line drain tubing is plugged.  Removed tubing from collection flask in refrigerator and cleared blockage with canned air.  RTS at 2052 GMT.

12/14/06

PAD

#1 sample line drain line is plugged.  There is residue visible inside collector flasks from sample pump that recently failed.  Removed drain line and found plug in adapter between drain line and flasks.  Removed collection flask to allow cleaning and replaced with spare flask found in PGS shed.  RTS at 2103 GMT.

11/24/06

PAD

Reference cylinder for PGS system is down to 300 psi.  Mentor request cylinder R/R.  Shut off reference cylinder, (XC003852B, 367 ppm.), at 2022 GMT.  Installed cylinder, (SG9164659BAL, 368.7 ppm.) and checked cylinder/ regulator connection for leaks.  Purged reference cylinder line for 2 minutes.  Adjusted low side of regulator to 20 psi.  New cylinder has 1960 psi.  RTS at 2042 GMT.

10/16/06

PAD

Instrument mentor has noticed 2m PGS pump flow is falling.  Observed Pressr_1 = 10.45 at 2012 GMT.  Shut down PGS pumps and R/R *2m pump.  Restarted PGS pumps.  Adjusted Pressr_1 flow to 13.5 psi.  RTS at 2042 GMT.   *Old - 1/640086   *New - 1/593123

10/11/06

CAM

West booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1505 GMT.  Raised booms and RTS at 2117 GMT.

9/19/06

PAD

Flow for sample line #1 and #3 is low and need adjusted.  Adjusted #1 sample flow for Pressr_1 of 11.1 to 12.5 and Pressr_3 of 12.1 to 12.9 at 1603 GMT.

9/19/06

PAD

#1 sample line drain tube is plugged.  Removed tubing and cleared obstruction with canned air.  RTS at 1613 GMT.

9/11/06

PAD

PGS computer clock is fast.  Opened Atomic Clock Sync program and pinged to set clock.  Computer clock was adjusted by 1 min. 40 sec.  RTS at 1936 GMT.

9/7/06

PAD

PGS pumps for 2m, 4m, 25m, and 60m needs to be rebuilt every six months.  Sebastien Biraud request pump rack assembly R/R with rack assembly previously sent to CF. Shut pumps 1-4 down at 1928 GMT.  R/R pump rack assembly.  Removed pump SNs 1-4 respectively:  1/593122, 1/593124, 1/640085, and 1/713615.  New pump SNs 1-4 respectively:  1/681968, 1/713616, 1/713617, and 1/593121.   Pumps restarted and system RTS at 1943 GMT.

8/29/06

PAD

#3 sample line drain is plugged.  Flask inside refrigerator is full and observed water in #3 flowmeter.  At 1850 GMT, removed sample 3 drain line and cleared obstruction.  Removed inlet and outlet lines from flowmeter and dried with canned air.  Adjusted sample pressures as follows:  Pressr_1  16.4 adjusted to 13.5,  Pressr_2  12.6 adjusted to 13.5. Pressr_3  18.77 adjusted to 13.5, Pressr_4  9.5 adjusted to 13.5.  RTS at 1905 GMT.

8/17/06

PAD

Reference cylinder is down to 310 psi.  Mentor requests cylinder changed.  At 1834 GMT, shut off cylinder (CC19266, 367.3 ppm), and installed cylinder (XC003852B, 367.0 ppm).  Checked for leaks and purged reference flow line for 1 min.  RTS at 1854 GMT.

8/3/06

PAD

Desiccant for PGS IRGA needs changed.  R/R desiccant at 1458 GMT.  RTS at 1513 GMT.

7/20/06

PAD

Sebastien Biraud requests the inlet line for the NOAA sampler moved from the inlet line for the CO instrument to the 60m inlet for the PGS system.  At 1912 GMT, disconnected inlet for the NOAA and capped off the tee.  Removed cap at the 4-way cross on the 60m inlet for the PGS and attached inlet for NOAA sampler.  RTS at 1932 GMT.

6/15/06

PAD

#1 sample flow value is 11.7 on Pressr_1.  Adjusted #1 sample flow to 14.0 at 1506 GMT.  RTS at 1521 GMT.

6/6/06

PAD

During PM rounds, observed #3 sample flowmeter has water in meter and indicator ball is stuck at 0.  #3 sample drain tube is plugged and collection flasks in refrigerator is full of water.  At 2045 GMT, removed #3 drain tube and cleared obstruction.  Removed #3 sample flowmeter inlet and outlet tube and dried with canned air.  Reattached tubes.  Flowmeter now reads .4 l/ min.  RTS at 2100 GMT.

5/30/06

PAD

#2 sample line discharge tube is plugged.  Removed #2 discharge line.  Found and removed blockage in connector at condensation collection flasks in refrigerator.  RTS at 1535 GMT.

5/30/06

PAD

PGS computer will not connect with network resources.  Allied Telesyn media converters in PGS shed and AOS trailer are not linked.  Reset media converters.  RTS at 1507 GMT.

5/25/06

PAD

PGS computer will not connect with network.  Reset Allied Telesyn media converters in PGS and AOS trailers.  RTS at 2020 GMT.

5/23/06

PAD

#1 sample flow drain line is plugged.  #1 sample flowmeter shows 0 l/ min.  After tapping on flowmeter, it registers .7 l/ min.  Removed #1 drain tube and cleared obstruction.  Removed inlet and outlet line from flowmeter and dried with canned air.  Flowmeter now shows .7 l/ min.  RTS at 1515 GMT.

5/9/06

PAD

#1 sample flow meter registers .3 l/ min.  Removed inlet and outlet tubes from flow meter.  A couple of drops of water came out of inlet tube.  Flushed flow meter with canned air and reattached tubes.  Flow meter now registers .7 l/ min.  RTS at 1554 GMT.

5/8/06

MK

While in PGS shed, observed #3 sample line drain tube plugged.  At 1900 GMT, removed tube and cleared obstruction with canned air.  RTS at 1915 GMT.

5/8/06

PAD

Instrument mentor request PGS reference cylinder changed.  At 1850 GMT, Mark Klassen and I shut off reference cylinder (XC003390B, 290 psi.).  Installed cylinder (CC19266, 2020 psi, 367.3 ppm).  Checked regulator connections for leaks.  Purged reference gas for 30 seconds.  RTS at 1905 GMT.

5/2/06

PAD

#1 sample flow meter is showing .3 l/ min.  At 1548 GMT, removed inlet and outlet lines from flow meter.  A couple drops of water came out of inlet line.  Dried flow meter with canned air and reconnected lines.  Flow meter now shows .7 l/ min.  RTS at 1603 GMT.

5/2/06

PAD

#3 sample line drain tube is plugged.  At 1540 GMT, removed drain tube and cleared obstruction.  RTS at 1555 GMT.

4/3/06

PAD

Notified by mentor he can access PGS computer remotely but is unable to transfer data files.  Mentor requests PGS computer rebooted.  Rebooted computer at 1928 GMT.  RTS at 1943 GMT.

3/21/06

PAD

Notified by Windows administrator that he could access PGS computer but was unable to move files.  Requested PGS computer rebooted.  Rebooted PGS computer at 2204 GMT.  Notified Windows administrator.  RTS at 2219 GMT.

3/21/06

PAD

# 3 sample drain line is not flowing.  Removed drain line at 1639 GMT and found dirt/ debris in connector between collection flask and drain line.  Removed blockage and assembled drain line.  RTS at 1654 GMT.

2/28/06

MK

The PGS computer is unable to communicate and collect data from the CR10X currently running the PGS system. Using a laptop computer, attempted to directly connect to CR10X.  No change. Power cycled CR10X, no change. Replaced 9-pin ribbon cable and tried again, still no change. R/R *CR10X and loaded PGS program. All communications were restored.   *Old – 32176   *New – 26909

2/14/06

PAD

Mentor request *PGS pumps changed every 6 months.  Removed rack containing old pumps #1-4.  Installed pump rack containing new pumps # 1-4.  Started pumps at 2133 GMT.  RTS at 2140 GMT.  

*Old pumps (#1-4) - 1/681968  1/713616  1/713617 1/593121

*New pumps (#1-4)  - 1/593122  1/593124  1/640085  1/713615

2/9/06

MK

Desiccant canister needs to be changed in PGS system. R/R desiccant canister at 1802 GMT. 

1/31/06

PAD

Sebastien Biraud request reference CO2 cylinder replaced in PGS shed.  At 1952 GMT, shut off reference cylinder ALM046365 (371.8 ppm CO2) and installed cylinder XC003390B (374.5 ppm CO2).  Checked regulator/ cylinder connection for leaks and purged line for 2 min.  New cylinder pressure is 1950 psi.

1/30/06

PAD

Sebastien Biraud notified #3 sample flow is low.  Checked for leaks around PGS rack and refrigerator.  No leaks were detected.  Adjusted Press_3 flow from 11.6 l/min to 12.5 l/ min.  Adjusted Press_4 flow from 15.0 l/ min to 14.0 l/ min. RTS at 2004 GMT.

12/7/05

PAD

Sebastien Biraud notified he couldn’t remotely access PGS, CO2met, CO2flux, or CO computers in PGS shed.     Reset Allied Telesyn media converters located in PGS shed and AOS trailer.  Computers can now access network.  Notified instrument mentor.  RTS at 1735 GMT.

10/21/05

PAD

PGS computer clock is 1 minute 45 seconds fast.  Atomic clock last successful ping was on 7/6/05.  Pinged atomic clock and synchronized clock at 1611 GMT.  RTS at 1626 GMT.

10/21/05

PAD

Reference cylinder pressure is down to 300 psi.  Mentor request new cylinder installed.  At 1601, removed cylinder number XC015008B, 370.8 ppm CO2, and installed cylinder number ALM046365, 371.8 ppm. CO2.  New cylinder pressure is 1930 psi.  Purged line and checked for leaks.  RTS t 1621 GMT.

10/20/05

PAD

# 3 sample line drain tube is plugged.  Removed drain tube and cleared obstruction with canned air.  RTS at 1540 GMT.

9/26/05

PAD

CO2met computer was found down.  Previous troubleshooting has found computer fan is not working.  Computer will shut down periodically.  Replacement fan is not available.  Mentor is aware of problem.  Restarted CO2met computer at 1416 GMT.  RTS at 1501 GMT.

9/26/05

PAD

# 3 sample line drain line is plugged.  Removed line and cleared obstruction.  RTS at 1435 GMT.

9/20/05

PAD

#3 sample drain line was plugged.  Removed and cleaned drain line at 1540 GMT.  RTS at 1555 GMT.

9/12/05

PAD

CO2met computer is off.  Started computer at 1534 GMT.  Computer fan is not running.  Notified instrument mentor.  RTS at 1549 GMT.

9/7/05

PAD

PGS system has had dry air running through system to dry out condensation since 9/1/05.  Mentor requests dry air removed and system returned to service.  Shut off dry air flow.  Removed condensation in sample lines from condensation tubes to outside of refrigerator with dry air.  R/R capillary drain tubes from condensation tubes.  Let refrigerator cool to below 10 deg. C.  Reconnected samples lines 1-4 to outlet lines from refrigerator.  Adjusted sample flows to approximately 12 psi.  Monitored values shown for system.  All values appear to be normal.  Notified instrument mentor.  RTS at 1730 GMT.

9/6/05

PAD

Facilities request all computers, instruments, UPS in PGS shed be shut down so repairs can be made to conduit located between PGS shed and 60m tower.  At 1753 GMT, shut down PGS, COflux, COmet, and CO computers.  Shut down all instruments and UPS located in PGS shed.  Restarted all computers, UPS, and instruments at 1837 GMT.  PGS computer will not connect with data logger and CO2flux computer will not connect with 25m and 60m IRGA and Sonic.  Further troubleshooting indicated 25m and 60m IRGA and Sonic were operating.  Rebooted computer and power cycled port box.  PGS computer was not detecting USB hardware cable coming from data logger.  Rebooted PGS computer and it connected with data logger.  RTS at 2033 GMT. 

9/6/05

PAD

During PM rounds, found CO2met computer was off.  Started CO2met computer.  Previous troubleshooting has found no problem with UPS or computer battery.  Checked power saver settings on computer.  No problems found with settings.  RTS at 1425 GMT.

9/1/05

PAD

Instrument mentor request dry air run through lines to attempt to dry out condensation in PGS system.  As instructed by mentor, at 1800 GMT, Chris Martin and I disconnected outlet lines 1,2,3, and 4 on refrigerator.  Connected dry air cylinder, supplied by mentor, to all 4 lines.  Set pressure of lines on computer display to 10 psi.  Removed lines from inline buffer cylinders to verify flow and check for condensation.  Flow was observed and no condensation was found.  Notified instrument mentor.  RTS at 1840 GMT.

8/26/05

PAD

PGS system has condensation in lines.  Mentor request inline buffer tubes removed and checked for water.  Removed buffer tubes and found approximately 2-3 cc of water in # 1 tube.  Drained water from tube and dried with canned air.  Notified instrument mentor.  RTS at 1925 GMT.

8/24/05

PAD

Instrument mentor request desiccant changed in PGS system.  Removed desiccant canister and replaced desiccant at 2037 GMT.  RTS at 2052 GMT.

8/24/05

PAD

PGS computers will not connect with network resources.  Found Allied Telesyn model AT-FH708E 8 port fast ethernet hub located in PGS shed is on raw power.  Plugged hub into UPS power.  Placed media converters located in PGS shed and AOS trailer into link/ test mode.  Communications were reestablished.  RTS at 1435 GMT.

8/23/05

PAD

Instrument mentor notified that PGS sample flow had dropped near 0 sometime around 2300 GMT on 8/22/05.  At 1900 GMT, went to PGS shed and called mentor.  Sample flow on computer display listed as MasPl_ccm is sporadic and ranges from -9999 to 100.  Mentor requested instrument rack power cycled.  Power cycled Isotope control box, valve box, and data logger box.  Flow value on computer is still sporadic.  Troubleshooting at this time.  RTS at 1915 GMT.

8/19/05

PAD

PGS sample line pumps were R/R yesterday.  Mentor request pressure on sample lines adjusted to 13 psi.  Adjusted to 13 psi.

8/18/05

PAD

Instrument mentor request sample line pumps for 2,4,25, and 60m replaced with pumps sent by mentor.  Shut down pumps at 1900 GMT.  Removed pump rack and brought to lab to replace pumps.  Replaced 2m pump SN1/593122 with SN 1/681968.  Replaced 4m pump SN 1/593124 with SN 1/713616.  Replaced 25m pump SN 1/640085 with SN 1/713617.  Replaced 60m pump SN 1/713615 with SN 1/593121.  Put new swagelock connectors on all inlet and outlet tubes and reinstalled pumps.  Restarted pumps at 2106 GMT.  RTS at 2121 GMT.

8/18/05

PAD

Instrument mentor informed he couldn’t remotely access computers in PGS shed.  Put Allied Telesyn media converters located in PGS shed and AOS trailer into link/ test mode.  Communications were restored.  RTS at 1430 GMT.

8/15/05

PAD

Mentor notified he cannot remotely access PGS, CO2met, or CO2flux computers in PGS shed.  PGS computer is running normally.  Put Ethernet media converters located in PGS shed and AOS trailer in link/test mode.  Communication was reestablished.  Computers will now connect with network resources.  RTS at 1948 GMT.  

8/12/05

PAD

Mentor requested 4m pump shut down and new pump installed with inlet line left disconnected.  Shut down pump #2, 4m, at 2117 GMT.  Installed new pump and connected outlet line.  Inlet line was intentionally left disconnected as per mentor's instruction until further notice.  Pump was started at 2125 GMT.  RTS at 2142 GMT.

8/11/05

PAD

Mentor notified he cannot remotely access PGS, CO2met, or CO2flux computers in PGS shed.  Computers are running normally.  Put ethernet media converters located in PGS shed and AOS trailer in link/ test mode.  Communication was reestablished.  Computers will now connect with network resources.  RTS at 1630 GMT.  

8/9/05

PAD

#1 sample line drain tube is blocked.  Flow meter has water in it. Removed drain tube and cleared obstruction with canned air.  Removed inlet line from flow meter and let water drain out of meter and dried with canned air.  RTS at 1656 GMT.

8/8/04

PAD

PGS computer will not connect with datalogger in PGS shed.  Checked                                                                   connections at computer and datalogger.  Power cycled datalogger.  Computer  would not connect.  Rebooted computer and it will now connect with datalogger.  RTS at 1919 GMT.

8/5/05

CAM

Need to power the PGS sample line pumps off for tower maintenance.  Powered pumps 1-4 off at 1328 GMT.  Powered pumps back on following maintenance at 1652 GMT.

8/4/05

CAM

Need to power the PGS sample line pumps off for tower maintenance.  Powered pumps 1-4 off at 1325 GMT.  Powered pumps back on following maintenance at 2035 GMT.

7/26/05

PAD

Sample line #1 drain tube is plugged.  Sample line flow meter has water in meter.  Removed sample line drain tubing and cleared line.  Removed inlet and outlet tubing from flow meter and removed water with canned air.  RTS at 1600 GMT.

7/13/05

PAD

Instrument mentor requests CO2 reference cylinder changed in PGS shed.  Removed cylinder CC192441, and installed cylinder XCO15008B.  Purged line and checked cylinder/ regulator connection for leaks.  Notified instrument mentor.  RTS at 1929 GMT.

7/6/05

PAD

#1 sample flow discharge tube from refrigerator is plugged.  Removed line and cleared obstruction at 1418 GMT.  Ball in sample line flowmeter was stuck.  Removed flowmeter and freed indicator ball.  Reassembled line and flowmeter.  RTS at 1438 GMT.

6/8/05

PAD

Instrument mentor notified 25m inlet pressure is not stable on computer display.  At 1817 GMT, 25m pressure 3_mV= -39.399.  Normal reading is about 15.0mV.  Disconnected/ reconnected pressure sensor.  Reading returned to normal.  Pressure3_mV= 15.738.  Notified instrument mentor.  RTS at 1832 GMT.

5/25/05

MR

The maintenance on the PGS equipment is finished and the SE elevators need to be raised.  Raised the SE elevators.  RTS at 2145 GMT.

5/24/05

MR

The SE elevators need to be lowered due to maintenance on the PGS equipment.  Lowered the SE elevators at 1410 GMT.  These will be left down overnight due to maintenance issues on the PGS equipment.

5/17/05

PAD

Reference cell flow observed at 3 mL/ min.  Adjusted reference cell flow to 8 mL/ min. at 1449 GMT.  RTS at 1504 GMT.

5/12/05

PAD

Reference cell flow observed at 3 mL/ min.  At 1455 GMT, adjusted reference cell flow to 8 mL/ min.  RTS at 1510 GMT.

5/10/05

PAD

At 1843 GMT, observed reference cell flow at 3 mL/ min.  Adjusted flow to 8 mL/ min. at 1845 GMT.  RTS at 1900 GMT.

5/3/05

PAD

#3 sampling line discharge tube from refrigerator is plugged.  This is the seventh time since 3/8/05 that line has been blocked.  At 1515 GMT, removed line and tried to clear blockage with can air and compressed air.  Blockage will not clear.  Discharge tubing is 30 cm in length.  I cut 12 cm off end of tubing.  Tubing will now flow air unrestricted.  RTS at 1615 GMT.

4/21/05

PAD

#3 and #4 sample line discharge tubes are blocked.  Removed lines and cleared obstruction with compressed air.  RTS at 1525 GMT.

4/5/05

PAD

#1 and #3 sample line discharge tubes are plugged.  Removed tubes and cleared obstruction.  RTS at 1441 GMT.

4/5/05

PAD

Facilities requested instruments shut down that receive power from breaker #5 in PGS shed so additional electrical outlets can be installed.  Shut down IRGA rack, vacuum pump, sample pumps, refrigerator, and unplugged modem adapter at 1345 GMT.  Restarted system at 1420 GMT.

3/31/05

PAD

Instrument mentor requests reference cylinder changed in PGS shed.  Removed cylinder # CC151933, 375.133 ppm CO2 and replaced with cylinder # CC192441, 370 ppm CO2.  Checked for leaks and purged line.  RTS at 2044 GMT.

3/29/05

PAD

#3 sample line drain tube is plugged.  At 1622 GMT, removed drain tube and cleared obstruction.  RTS at 1637 GMT.

3/24/05

PAD

Reference cell flow observed at 3 mL/ min.  Adjusted reference cell flow to 8 mL/ min at 2004 GMT.  RTS at 2019 GMT.

3/24/05

TMG

Stopped PC208W software @ 1830 GMT and installed RAdmin @ 1840 GMT. Rebooted the computer @ 1844 GMT and RTS @ 1848 GMT.   Tried to install RAdmin on the CO2Met but was not successful.

3/22/05

PAD

#3 sample line drain tube is plugged.  Removed tube and cleared obstruction.  RTS at 1715 GMT.

3/22/05

PAD

Reference cell flow observed at 3 mL/ min.  Adjusted reference cell flow to 8 mL/ min.  RTS at 1709 GMT.

3/8/05

PAD

#2 and #3 sample line discharge tubes are plugged.  Removed discharge lines and cleared obstruction at 1845 GMT.  RTS at 1900 GMT.

3/3/05

PAD

#1 sample line discharge tube from refrigerator is plugged.  Removed tubing and cleared obstruction.  RTS at 1650 GMT.

3/1/05

PAD

#3 sample line discharge tube from refrigerator is plugged.  Removed tubing and cleared obstruction.  RTS at 2000 GMT.

2/24/05

PAD

Reference cell flow observed at 3 mL/ min.  Adjusted reference cell flow to 8 mL/ min. at 2018 GMT.  RTS at 2033 GMT.

2/22/05

PAD

Reference cell flow observed at 3 mL/ min.  Adjusted reference cell flow to 8 mL/ min. at 2015 GMT.  RTS at 2030 GMT.

2/22/05

PAD

#1 sample line drain tube coming from refrigerator is plugged.  Removed tubing and cleared obstruction.  Reconnected tubing and checked for flow.  RTS at 2025 GMT. 

2/22/05

PAD

At 1718 GMT, found CO2 flux and CO2 met computers were off.  PGS sample pumps were not running.  Computer monitor did not have power.  Found tripped breaker in PGS shed caused by shorted wire in PGS pump rack.  Reset breaker and replaced blown fuse in pump rack.    Removed pump rack assembly and removed pump #2.  Discovered shorted wire that had rubbed through insulation against #2 pump.  Repaired wire and installed #2 pump.  Installed pump rack assembly and restarted pumps at 2003 GMT.  RTS at 2018 GMT. 

2/15/05

PAD

#2 sample line discharge tube from refrigerator is plugged.  Removed tubing and cleared obstruction.  RTS at 1825 GMT.

2/11/05

PAD

Instrument mentor requested Envirogen desiccant in PGS IRGA system be changed out with Dri-Rite desiccant.  Changed desiccant at 2007 GMT.  Notified instrument mentor.  RTS at 2022 GMT.

2/8/05

PAD

#3 sample line discharge tube from refrigerator is plugged.  Removed line and cleared obstruction.  RTS at 1625 GMT.

1/27/05

PAD

Reference cell flow observed at 4 ml/ min.  Adjusted reference cell flow to 8 ml/ min at 1610 GMT.  RTS at 1625 GMT.

1/25/05

PAD

#1 and #3 sample flow line discharge tubes from refrigerator are plugged.  Disconnected lines and tried to blow through lines to remove blockage.  Could not get #1 line clean.  Removed both lines and cleared obstruction with canned air.  Replaced #1 and #3 discharge lines and checked for flow.  RTS at 1735 GMT.

1/20/05

PAD

#1 and #2 sample drain tubes from condensation flasks in refrigerator are plugged.  Removed tubes and cleared obstruction.  RTS at 1954 GMT. 

1/11/05

PAD

Reference cell flow is 4 ml/ min.  Adjusted reference cell flow to 8 ml/ min. at 1618 GMT.  RTS at 1633 GMT.

1/7/05

PAD

Pump #4 was shut down due to ice blocking inlet tube.  Started pump #4 at 2010 GMT.  #4 sample flow shows .4 l/m.  Computer shows Pressr_4  12.372.   Notified instrument mentor.  RTS at 2025 GMT.

1/6/05

PAD

Sebastien Biraud informed sample flow #4 from 60m level on tower was low.  Found sample flow was not registering on computer.   Suspect line plugged with ice.  Shut off pump #4 at 2055 GMT.  Notified instrument mentor.

1/6/05

PAD

Reference cell flow observed at 3 ml/ min.  Adjusted reference cell flow to 9 ml/ min. at 1545 GMT.  RTS at 1600 GMT.

12/30/04

PAD

Reference cell flow observed at 4 ml/ min.  Adjusted reference cell flow to 8 ml/ min. at 1621 GMT.  RTS at 1636 GMT.

12/28/04

PAD

During PM rounds, #1 drying tube from collection flask located in refrigerator was found plugged.  Removed line and drained collected moisture from collection flask.  Cleared obstruction from tube and reassembled.  RTS at 2233 GMT.

12/23/04

PAD

Reference cell flow meter shows 3 mL/ min. and is out of range.  Adjusted reference cell flow to 9.0 mL/ min. at 2015 GMT.  RTS at 2030 GMT.

12/21/04

PAD

Instrument mentor requested CO2 reference cylinder bottle replaced.  Shut off cylinder SN CC185567, (372.8 ppm.) at 1624 GMT.  Installed cylinder SN CC151933, (375.133 ppm.).  Checked for leaks at cylinder and regulator.  Purged reference flow line.  New cylinder shows 1910 psi.  RTS at 1644 GMT.

12/21/04

PAD

Reference cell flow meter observed at 3 ml/ min.  Increased flow to 8 ml/ min.  RTS at 1630 GMT.

12/7/04

PAD

Reference flow rate observed as 4 ml/ min.  Adjusted flow rate to 10 ml/ min. at 1638 GMT.  RTS at 1653 GMT.

11/30/04

PAD

Sebastien Biraud notified he could not remotely access PGS computer.  Computer appears to be going into standby mode.  Checked power options in computer control panel.  Power scheme is Dell low power setting.  Changed setting to Always on.  Computer was set to go to system standby after 30 min.  Changed setting to never.  RTS at 1608 GMT.

11/29/04

PAD

Sebastien Biraud notified he could not access PGS computer.  PGS computer power light was flashing.  Pushed power button on computer and logged on.  Notified instrument mentor.  RTS at 2105 GMT.

11/29/04

PAD

Sebastien Biraud notified he could not access PGS computer.  PGS computer power light was flashing.  Pushed power button on computer and logged on.  Notified instrument mentor.  RTS at 1922 GMT.

11/16/04

PAD

Sampling drain tube #1 from refrigerator is plugged.  #1 collector in refrigerator has condensation.  Removed and cleared obstruction from drain line #1.  Sample line flow dropped from 2.0 l/m to 1.3 l/m.  RTS at 1902 GMT.

11/12/04

PAD

Sebastien Biraud emailed he cannot remotely access PGS computer.  Email was received on 11/11/04, which is a holiday.  Power cycled network hub.  Computer will now access network resources.  Notified instrument mentor.  RTS at 1550 GMT.

11/8/04

PAD

Sebastien Biraud notified he cannot remotely access PGS or CO2met computers.   Power cycled network hub and fiber optic modem.  Computers will now access network resources.  Notified instrument mentor.  RTS at 1751 GMT.

10/26/04

PAD

At 1945 GMT, #1 sample flowmeter was observed stuck at 5.0 l/m and had water in flowmeter.  #1 and #2 sample flow drain lines from tubes in refrigerator were plugged.  Removed #1 and #2 drain lines and cleared obstructions.  Removed inlet line from sample #1 flowmeter and drained water from flowmeter.  RTS at 2015 GMT. 

10/20/04

MR

The instrument mentor requests that the calibration on the 60M IRGA be performed again.  Lowered the tower at 1750 and removed the IRGA.  Performed the calibration.  Raised the tower at 1930 GMT.  Notified the mentor.

10/12/04

PAD

Instrument mentor requested standard reference cylinders replaced.  Dave Billesbach is on site and I assisted him.  Removed cylinders CA04434 (335ppm.), CA04437 (360ppm.), CA04404 (390ppm.), CA0444 (420ppm.).  Installed cylinders CA05655 (355ppm.), CA05687 (380ppm.), CA05679 (420ppm.), CA05658 (460ppm.).  Purged lines from cylinders to rack and checked for leaks.  Dave Billesbach was doing calibration checks after system was RTS at 2030 GMT.

10/12/04

PAD

Sebastien Biraud notified he could not access PGS computer.  Power cycled network hub and fiber modem adapter at 1447 GMT.  Computer will now connect with network resources.  RTS at 1502 GMT.

10/8/04

PAD

Site had a power bump at 1605 GMT.  Checked PGS system.  All Sonic programs were running but computer could not access network resources.                                                    Power cycled network hub and fiber optic modem.  Made sure computer would access network resources.  RTS at 2015 GMT.

10/7/04

PAD

PGS computer locked at 1939 GMT.  Logged off/on as operator at 1947 GMT.  RTS at 2002 GMT.

10/7/04

PAD

#2 sample drying line from refrigerator is plugged.  Removed #2 line from collection flask inside refrigerator and removed obstruction.  RTS at 1915 GMT.

9/27/04

CAM

West booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1432 GMT.  Raised booms and RTS at 2105 GMT.

9/9/04

PAD

During PM rounds, sample flow #1 tube was found plugged.    At 1558 GMT, removed tubing from #1 collection flasks located in refrigerator.  Cleared blockage and made sure tube was cleared.  RTS at 1613 GMT.

8/31/04

PAD

Instrument mentor requested reference cylinder changed.  Shut off reference regulator at 1908 GMT.  Removed cylinder # CC91963 (373.6 ppm.) and replaced with cylinder #CC185567 (372.8 ppm).  Had to try three different regulator connectors before finding one that would not leak.  Purged reference line and adjusted cylinder regulator to 22 psi.  Secondary reference regulator is set at 15 psi.  RTS at 1948 GMT.

8/30/04

PAD

Sebastien Biraud informed that IRGA flow for channel 1 was low.  Requested system checked.  Checked system for leaks and checked water trap tubes in fridge to make sure they had flow.  Adjusted channel 1 flow from .4 to .7 on sample flow meter.  Pressr_1 reads 12.972 and Press1_mV reads 16.775.  Notified instrument mentor.  RTS at 1836 GMT.

8/13/04

PAD

Sebastien Biraud reported Channel 1 pressure has shown a lot of fluctuation the past week and asked that connections be checked.  At 2027 GMT, Mike Rainwater and I checked connections for Channel 1.  Tightened connections at inlet block located on wall behind rack and buffer cylinder on table.  Computer displays Pressr_1  11.8 and Pressr_1mV 15.1.  RTS at 2042 GMT.

8/12/04

PAD

Sample gas flow #4 is barely registering on flow meter.  On PGS computer, #4 values for Pressr_4 is 8.9 and Press4_mV is 11.5.  Values are lower than other sample flows.  At 1924 GMT, adjusted #4 sample flow to read Press4_ 12.0 and Press4_mV 15.5.  Values are now in line with other sample flows.  RTS at 1939 GMT.

8/11/04

PAD

PGS and CO2met computers will not access network resources.  Power cycled network hub and fiber optic modem.  PGS and CO2 met computers will now access network resources.  Power cycled camera in PGS shed.   CO2met computer will now communicate with camera.  Notified instrument mentor.  RTS at 1604 GMT.

8/5/04

MR

The SW elevator needs to be lowered due to SN# need to be checked.  Lowered the tower at 1450 GMT.  Raised the tower at 1523.  RTS at 1530 GMT.

7/23/04

PAD

Instrument mentor cannot remotely access PGS or CO2 flux computer.  Power cycled network hub and fiber optic modem.  Computer will now access network resources.  Notified instrument mentor.  RTS at 1848 GMT.

6/30/04

MR

The power to the site was lost at 1330 GMT.  Power was restored to the site at 1455.  Powered up the computer system at 1455 GMT.  System came up and rebooted ok.  RTS at 1510 GMT.

6/30/04

MR

The new telebyte fiber mux has arrived on site and needs to be installed. Installed the new fiber mux in the fiber box at the 4M CO2 Flux system.  Tested communications and they tested good.  RTS at 1525 GMT.    *Old - LL2346   *New - QQ4699

6/29/04

PAD

Instrument mentor notified sample gas flows # 2 and 4 be adjusted.  Sample values on computer display are below other sample flow values.  Adjusted sample # 2 and 4 at 1454 GMT.  Values are now in line with other sample flows.  Notified instrument  mentor RTS at 1456 GMT.

6/24/04

PAD

Instrument mentor requested #2 sample gas flow adjusted.  Adjusted sample gas flow #2 in PGS rack from 1.7 to .8 at 1532 GMT.  Display shows Pressr 2= 12.198 and Press2 mV= 15.910.  Values are now in line with other sample flows.  RTS at 1547 GMT.

6/17/04

PAD

#2 sample gas flow meter was reading 3.0.  Collection tube in refrigerator was full of water.  At 1520 GMT, removed drain line from #2 collection tube.  Line was open.  Removed connector from collection tube and cleared obstruction.  Cleaned connection and replaced drain line.  RTS at 1535 GMT.

6/11/04

PAD

Instrument mentor requested more troubleshooting done on communication problem between 4m data logger and CO2met computer.  Chris Martin and I stopped Sonic 1 at 1718 GMT.  Removed covers of Telebyte fiber mux boxes in PGS shed and 4m tower.  Checked to make sure jumper pins were on 3 and 4.  Replaced boxes and switched data logger to port 2 of mux and downloaded data.  Synchronized data logger clock with computer after download.  Data logger was 4 min. 53 sec. slow.   Restarted Sonic 1 at 1815 GMT.   Contacted instrument mentor and informed of results.   RTS at 1818 GMT.

6/8/04

PAD

Troubleshooting ongoing communication problem with data logger at 4m tower and CO2Met computer.  Instrument mentor requested ground wire checked between computer and communications box in PGS shed.  At 1358 GMT, Chris Martin and I checked ground wire.  We power cycled fiber mux hubs at 4m tower and in PGS shed.  Switched data logger cables from port 3 to port 2 that is working.  Data logger will now connect with computer.  Switched cables back and notified instrument mentor of results.  Instrument mentor is troubleshooting at this time.  Waiting on reply.  RTS at 1438 GMT.

6/7/04

PAD

Igor Pensenson emailed that CO2flux system had gone down sometime the previous day and he could not remotely access PGS computer.  At 1352 GMT, all CO2 flux systems were graphing.  Computer cannot access network resources.  Power cycled network hub and modem changer at 1354 GMT.  Computer can now access network resources.  Contacted instrument mentor.  RTS at 1408 GMT.

6/3/04

PAD

During PM rounds, 4m CO2 met computer would not connect with data logger.  Severe weather at site yesterday.  Mike Rainwater and I hooked laptop to data logger and could communicate with laptop.  Power cycled port hubs located at base of 4m tower and inside PGS shed.  Troubleshooting at this time.  Instrument mentor has been notified.

5/20/04

PAD

PGS or CO2 flux computers cannot access network resources.  Power cycled network hub at 1624 GMT.  Computer will now access network resources.  RTS at 1639 GMT.

5/17/04

PAD

Sebastien Biraud notified he cannot remotely access PGS or CO2 flux computers.  Power cycled network hub at 1640 GMT.  Computer will now access network resources.  Notified mentor connection was restored.  RTS at 1655 GMT.

5/3/04

PAD

Sebastien Biraud informed he is unable to remotely communicate with PGS or CO2 flux computers.  Power cycled network hub at 1713 GMT.  Computer would still not access network.  Power cycled network hub and fiber optic modem.  Computer will access network.  Informed instrument mentor.  RTS at 1728 GMT.

4/29/04

PAD

Instrument mentor requested reference cylinder changed out.  Shut off cylinder valve at 1844 GMT.  Removed reference cylinder CC112427, (361ppm. CO2).  Installed cylinder CC91963, (373.6 ppm).    Fixed leak at secondary regulator.  Opened valve and bled air through bleed valve.  RTS at 1914 GMT.

4/15/04

CM/PD

SE booms need to be lowered for 25M MFR maintenance.  Powered PGS pumps off and lowered booms at 2104 GMT.  Raised booms and powered on PGS pumps at 2151 GMT.

2/12/04

MR

The 60M sonic is not communicating with the computer.  T/S found a bad RS 422 bit driver.  Cleaned the fiber connections on the bit driver.  Communications was restored.  RTS at 1715 GMT.

2/11/04

MR

The 60 M Sonic is not able to communicate with the computer. Lowered the elevator at 1614 GMT.  Connected the jumpers up to the system and found that the problem did not change.  Further T/S found that the elevator wiring is not correct.  Corrected the wiring on the 60M elevator.  Raised the elevator at 2230 GMT.

2/10/04

MR

Tower systems is finished with the tower and the elevators need to be returned to their operating position.  Raised the towers at 1800 GMT.  RTS at 1830 GMT.

2/9/04

MR

Tower systems is here and will be working on the 60M tower connections. Lowered the 25M and 60M elevators at 1600 GMT.  They will remain lowered until tower systems is finished with the work.

1/22/04

PAD

At 1930 GMT, pump #4 in PGS pump rack was observed locked up.  After collecting NOAA samples that draw air samples from line 4, shut off pumps on pump rack at 2035 GMT.  R/R *pump #4.  Turned on pumps at 2111 GMT and made sure all pumps came back normally.  Contacted instrument mentor and notified system returned to service.  Sent pump SN 1/593123 to instrument mentor for repairs.  RTS at 2115 GMT.  *Old - 1/593123  *New - 1/713615

1/14/04

PAD

Igor Pesenson e-mailed he cannot remotely communicate with CO2met or CO2 flux computers.  Power cycled network hub and fiber optic modem at 1658 GMT.  Computers will now access network.  Notified instrument mentor.  RTS at 1711 GMT.

1/9/04

CF

Instrument mentor requested reference cylinder changed in PGS shed.  At 2148 GMT, removed reference cylinder # CC149239 (373 ppm. CO2 concentration).  Installed cylinder # CC112427 (361 ppm. CO2 concentration).  Purged air from line.  Notified instrument mentor that cylinders had been changed out.  RTS at 2203 GMT.

12/11/03

PAD

Notified by instrument mentor pump #3 was showing erratic readings.  Unplugged and reseated connector on pump #3 pressure sensor at 1525 GMT.  Connected data logger on computer and made sure reading returned to normal.  Notified instrument mentor. RTS at 1535 GMT.

12/4/03

PAD

Rodents have been chewing heat flux wires that have to be spliced when broken.  Instrument mentor requested we place conduit over heat flux wires coming out of ground and suspend wires above ground to 4m tower.  At 1925 GMT, Mike Rainwater and I split 3 ft. pieces of conduit and placed heat flux wires inside.  Dug up heat flux plates, buried conduit with wires in them, brought up out of the ground and attached to stakes.  Suspended wires above ground to 4m tower.  Spliced #4 heat flux plate wire that had been chewed into by rodents.  RTS at 1955 GMT.

11/12/03

PAD

Igor Pesenson e-mailed he cannot remotely communicate with the PGS or CO2 Flux system computers.  Power cycled network hub at 1448 GMT.  Computer would not connect to any network resources.  Power cycled fiber optic modem.  Computer will now access network resources.  Notified instrument mentor connection was restored.  RTS at 1500 GMT.

10/23/03

PAD

Igor Pesenson informed he is unable to remotely communicate with the PGS computer.  Rebooted the network hub in the PGS shed at 1907 GMT.  PGS computer can now be accessed remotely.  Notified Igor access is restored.  RTS at 1920 GMT.

10/2/03

PAD

Reference cylinder regulator connector is leaking.  Replaced regulator connector with new one.  Tightened connector and checked for leaks.  It did.  Troubleshooting with instrument mentor at this time.

10/1/03

PAD

Heat flux plates for 4m CO2 flux system had been removed so farmer could plant wheat under instrument boom.  Buried heat flux plates 5 cm and checked computer to confirm values.  Heat flux 1 had been chewed by rodents and is not operational.   Notified mentor of installation.  RTS at 2015.

9/30/03

PAD

Weekly PM report states desiccant needs to be replaced.  Disconnected lines and removed old desiccant.  Filled drierite with new desiccant and reconnected lines.   RTS at 1947 GMT.

9/30/03

PAD

Connection between the reference cylinder and the regulator has been leaking for some time.  Troubleshooting to find cause.  Shut off reference cylinder at 1855 GMT.  Changed regulator connector on reference cylinder with connector from STD 420 regulator.  Leak on reference cylinder was worse.  Switched connectors back to original regulators.  Troubleshooting with AirGas at this time.  RTS at 1930 GMT.

9/25/03

PAD

Reference cylinder needed to be changed out at instrument mentor's request.  Shut off reference flow at 1929 GMT.  Removed old cylinder SG9147358, actual concentration of 362.5 ppm.  Replaced it with new cylinder CC149239, certified concentration 373 ppm.  Purged line and checked for leaks at regulator connection.  RTS at 1943 GMT. 

9/17/03

PAD

Soil probes for 4m CO2 flux system in field 8 needs to be removed so wheat can be planted under instruments on horizontal bar.  Removed soil moisture, soil temperature, and heat flux plates from under instrument horizontal bar.  3 soil moisture probes were not reading acceptable values prior to removal.  3 soil temperature probes ends came off when removing cables from soil trenches.  1 heat flux cable had been chewed by a rodent.  Will notify instrument mentor concerning removal time and damaged cables.  RTS at 2015 GMT.

8/25/03

JEM

Igor informed that he is unable to remotely communicate with the PGS or the CO2 Flux System.  The PGS computer is unable to access any network resources.  Rebooted the Network Hub in the PGS shed @1445 GMT.  The PGS computer is now able to access network resources.  Notified Igor that access to the PGS and CO2 Flux should be restored.  RTS @1455 GMT.

8/20/03

PAD

Instrument Mentor discovered 4m CO2 flux system had not collected slow data since 8/06/03.  Laptop would not connect with PC208W data logger.  Rebooted computer at 1600 GMT.  Checked that computer would connect with data logger and notified instrument mentor.  RTS at 1610 GMT.

8/15/03

PAD

Standard cylinder 390 secondary reading was .5.  Contacted instrument mentor.  Adjusted secondary flow to 12 at 2035 GMT.  RTS at 2050 GMT.

8/15/03

PAD

Sample gas flow meter 1 was reading 2.0.  Contacted instrument mentor and adjusted sample 1 flow to .6 at 2030 GMT.  RTS at 2035 GMT.

8/14/03-8/15/03

CAM

Dave Cook on site for annual 25 and 60M T/H calibration changeouts and temperature sensor ice baths.  Tower Systems also on site for tower inspection and upgrades to PGS filter systems.  West booms lowered at 1515 GMT on 8/14 for Tower Systems maintenance and air filter assembly upgrades.  The 60M level plunger connectors were discovered to have cracks developing on the boom portion of the connector.  To minimize stress, the booms were left down overnight.  Returned on 8/15 for sensor calibration changes.  Raised booms and RTS at 1939 GMT on 8/15.  

8/15/03

CAM

Need to lower SE booms for annual 25 and 60M T/H change outs.  Lowered the SE booms at 1341 GMT for maintenance.  Raised booms and RTS at 1920 GMT.

8/14/03

CM

Tower Systems on site.  SE booms lowered a few inches at 1900 GMT to tension the cable between the ground and 25M boom.  Raised booms at 1914 GMT.  Only the 60M boom made connection at that time due to the added tension.  Lowered booms again at 1925 GMT and repositioned the 25M boom higher on the cable.  Raised booms and RTS both 25 and 60M levels at 1935 GMT.

8/14/03

CM

Tower Systems on site.  SE booms lowered for tower system maintenance.  Sample flow pumps were shut down at 1443 GMT.   Lowered booms at 1443 GMT.  Replaced the 25 and 60M air filter assemblies.  Raised booms at 1515 GMT.  Pumps RTS at 1525 GMT.

8/05/03

PAD

CO2 flux system in field 21, data logger did not transfer data to laptop computer on slow side for download.  Found broken wire between data logger and laptop computer.  Disconnected wires at 1955 GMT and fixed broken wires in lab.  Re-installed wires to data logger and checked to make sure computer could connect to data logger.  RTS at 2040 GMT.

7/29/03

PAD

Soil sensors were pulled from the 4m tower so farmer could till ground under instruments.  Dug new trenches and reinstalled soil moisture probes, temperature sensors, and heat flux plates.  One of the heat flux cables had a broken wire, which was repaired.  Trenches had to be watered due to dry conditions to make good soil contact with moisture sensors.  Checked laptop data after installation to make sure sensors were working.  Soil moisture sensor #1 is not reading in acceptable range.  Notified mentor of installation and sensor readings.  RTS at 1650 GMT.

7/25/03

PAD

Soil moisture sensors, temperature sensors, and heat flux plates for the CO2 flux system needed to be taken up in order for the farmer to till the ground under the instruments at the 4m tower.  Removed soil moisture sensors, temperature sensors, and heat flux plates at 1440 GMT.  Spoke with tractor operator about instruments and informed him when sensors were removed.  Contacted instrument mentor and informed him of sensor removal.  RTS at 1840 GMT.

7/24/03

PAD

Upgrade communications and power lines from PGS shed to 4m tower located at base of 60m tower cable support.  Disconnected power and communication cables from 4m tower and PGS shed on 07/23/03 at 1410 GMT.  Contractor installed and buried 2- 1" conduits and pulled power and communication cables through buried conduits.  Power and communication cables were reconnected at 4m tower and PGS shed.  RTS on 07/24/03 at 1520 GMT.

7/3/03

MR

The Reference cylinder has used more gas then normal.  T/S found a leak in the line at the valve connection.  Tightened the connection.  RTS at 2007.

6/30/03

JEM

The reference cylinder pressure is low and needs to be changed out.  Disconnected the reference cylinder @1925 GMT.  Placed the new cylinder in service @1930 GMT.  The cylinder pressure was 200PSI.  The secondary pressure was set to 10PSI.  RTS @1935 GMT.

6/24/03

PAD

Reference cylinder secondary pressure was 19 psi.  Adjusted reference cylinder secondary pressure to 15 psi at 1450 GMT.  RTS at 1505 GMT.

5/16/03

PAD

Glass collection tube in Line 3 inside refrigerator was full of water at 1827 GMT.  Pulled drain line off, drained water out of collection tube and cleared drain line.  RTS @ 1842 GMT.

5/15/03

JEM

Sample flow meter #4 is full of water and the ball is floating at the top.  Disconnected the lines from sample flow meter #4 @1500 GMT.  Disconnected several different fittings along line #4 to allow the lines to drain out any water inside.  Blew all of the water out of the sample flow meter using the air compressor.  Reconnected all fittings on line #4 and reinstalled the flow meter @1520 GMT.  The flow then returned to 0.6.    RTS @1530 GMT.

5/15/03

JEM

Pump #3 has previously been installed on the wall rack and needs to be placed into the pump housing with the other pumps.  Powered off and disconnected all 4 pumps in the housing @1419 GMT.  Removed the bad pump #3 from the housing and installed the good pump.  Reconnected and powered on the 4 pumps @1447 GMT.  All pressures returned to normal levels.  RTS 1449 GMT.

5/1/03

JEM

The IRGA desiccant needs to be replaced because it is at 180mm.  R/R IRGA desiccant @1910 GMT.  RTS @1922 GMT.

4/15/03

JEM

Tubes 2, 3, and 4 in the drying sample refrigerator are not bubbling.  Removed the copper lines from drip flasks inside the refrigerator.  Each of the 3 lines was plugged with a small amount of debris.  Removed debris by blowing back through the lines.  Reinstalled the copper lines to the drip flasks.  Lines 2, 3, and 4 are now bubbling normally.  RTS @1452 GMT.

4/1/03

JEM

The IRGA Desiccant needs to be replaced.  R/R IRGA Desiccant @1527 GMT.  RTS @1535 GMT.

3/24/03

JEM

Have been notified by Igor Pesenson that the 4 pumps are showing zero flow and have been since Friday (3/21/03).  Upon checking the PGS, the pumps were found to be powered off.  Pumps where most likely left off by accident when the 60M tower was returned to service.  Powered the pumps on @1700 GMT.  RTS @1715 GMT.

3/17/03

JEM

Pumps 1,2,3 and 4 were shut down @1635 GMT while the Southeast booms were lowered on the 60M tower.  Nothing is currently wrong with the PGS.  RTS @1705 GMT.

3/13/03

JEM

Igor notified that the reference pressure is dropping faster than expected.  Checked all lines inside the PGS shelter for leaks.  Two leaks were found.  A small leak on pump #1 outlet side.  And a small leak at the head of the reference cylinder where the gauges connect.  Powered pump #1 off @2005 GMT to make repairs to a split found in the tubing.  Repaired the tube and reinstalled it @2015 GMT and checked for leaks  -  None were found.  Removed the gauges from the reference cylinder and blew the connector out to remove any dirt or debris that may be present.  Reinstalled the gauges and checked for leaks  -  None were found.  RTS @2025 GMT.

3/10/03

JEM

Notified by the mentor that line 3 has 0 pressure.  Upon checking the PGS, pump #3 was observed locked up.  Removed the lines from the pump and it still would not operate.  Powered off all 4 pumps @1737 GMT.  R/R Pump #3 @1745 GMT.  Powered on all 4 pumps @1755 GMT.  Pump #3 began working and built pressure up to 9.8 PSI, and then it also locked up.  Removed the lines from the pump, but it still would not operate.  R/R pump #3 again @1845 GMT.  At 1855 GMT, line #3 was reading 10.6 PSI.    RTS @1900 GMT.

3/3/03

JEM

The CO2 cylinder has become empty and needs to be replaced. FedEx arrived with the replacement cylinder.    Removed the empty cylinder @2203 GMT.  Installed the new cylinder @2207 GMT.  Opened the new CO2 cylinder @2210 GMT.  The pressure reading on the gauge was 2000psi.  RTS @2215 GMT.

2/21/03

MR

The tower pumps in the system need to be shut down  due to maintenance on other equipment.  Shut the pumps down at 1925 GMT.  Turned the pumps back on at 2115 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

1/30/03

JEM

The 60M Southeast temperature has been to high.  There is no problem with the PGS.  Corrective maintenance is being done on the 60M T/H.  Lowered Southeast 25M and 60M booms @1655 GMT.  RTS @1810 GMT.

1/16/03

JEM

The refrigerator temperature is -2.0 degrees Celsius.  Adjusted the thermostat on the refrigerator to bring the temperature above freezing.  RTS @1630 GMT.

1/13/03

MR

The instrument mentor wants the NOAA flask sampler removed and sent in for repairs.  Removed the NOAA flask sampler.  RTS at 1615 GMT.

12/16/02

MR

The tower needs to be lowered due to maintenance on other equipment.  Lowered the tower at 1639 GMT.  Raised the tower at 1705 GMT.  RTS at 1710 GMT.

11/25/02

MR

Line 4 has water inside the tube and in the glass separation tube.  T/S found this line was plugged.  Replaced the line with another of the same size and drained the water from the tube.  RTS at 1725 GMT.

11/14/02

KT

The desiccant needed to be changed out.  At 1727 GMT removed the desiccant housing.  At 1730 GMT removed and replaced the desiccant.  RTS at 1742 GMT.

11/7/02

CM/MR

SE booms lowered from 1557 GMT to 1755 GMT to troubleshoot 25M IRT communications problems.  When the booms were raised, lines 1 and 3 flow rates were observed to be near zero.  Reseated the booms and line 3 low returned to normal but line 1 remained zero.  Traced #1 air line out, disconnecting the unions along the way to find the source of blockage.  Found that there was flow up to the refrigerator, but none coming out.  Opened the ice box door for a minute or so after which time line 1 was observed to physically move as if an obstruction came free.  Rechecked line 1 air flow and it returned to 0.8 on the flow meter and 13.2 psi on the computer reading.  Suspect there may have been ice in the line.  RTS status O.K. at 2005 GMT with all flow rates normal.

11/6/02

CAM

SE booms lowered from 2017 GMT to 2222 GMT to troubleshoot 25M IRT communications problems.  The PGS pumps were powered off while the booms were down.

11/5/02

CAM

PC208 program would not plot graphs or display numerical data.    Rebooted computer at 2022 GMT and RTS status O.K.

11/1/02

CAM

Instrument mentor requested that pump #3 flow rate and pressure be checked.  At 1755 GMT the flow meter ball was practically at 0 l/min and barely moving at all.  The pressure was reading around 8 psi.  Checked all pump connections and found that both line 1 and line 3 were leaking on the pressure side of the pump.  Tightened connections and line 3 pressure increased slightly to around 8.5 psi.  Mentor requested that the external pump be reconnected to line 3 to see if conditions improve.  Connected the external pump and the pressure decreased slightly.  Mentor requested that another new pump be installed on line 3.  Powered all pumps off at 1904 GMT.  Replaced pump #3 with a new spare and powered pumps back on at 2052 GMT.  Rechecked pressure and observed no significant change.  At 2100 GMT lines 3 and 4 were swapped between pumps to isolate the cause of reduced pressure.  T/S revealed that the pressure loss is occurring in vacuum line #3 that runs up to the tower.  Returned all lines to their original locations and RTS at 2108 GMT.  Notified mentor of results.

10/31/02

CM/MR

SE booms lowered from 1445 to 1618 GMT for 25MFR and 60 T/H maintenance.

10/29/02

CAM

SE booms lowered from 2135 GMT to 2220 GMT for 25 and 60M T/H maintenance.

10/28/02

CAM

The SE 25 and 60M T/H sensors are reading –9999’s or 0’s.  Lowered the SE booms slightly at 1731 GMT and then raised to reseat connections.  RTS at 1733 GMT.

10/10/02

MR

The instrument mentor requested that the desiccant in the gas sample line be changed. Changed the desiccant.  RTS at 1945 GMT.

10/10/02

MR

The instrument mentor wants the CO2 tank changed. R/R the CO2 tank.  set the output regulator pressure to 10 PSI as requested.  RTS at 1930 GMT.

9/23/02

MR

The Southeast elevator needs to be lowered due to work being performed on the 25M MFRSR.  Lowered the tower at 1930 GMT.  Raised the tower at 1943 GMT.  RTS at 1945 GMT.

9/3/02

MR

The instrument mentor wants the 60M level elevator lowered and the air sampling tubes checked to make sure there is no blockage.  Lowered the tower.  The connection seemed OK.  Changed the filter due to it looking dirty.  This did not remove the blockage.  Suspect the blockage is in the connector that is attached to the tower.  RTS at 2040 GMT.

8/21/02

MR

Pump # 3 has locked up again and needs to be replaced.  Called the mentor and she agreed that pump # 3 should be replaced.  The spare pumps are too small for a direct replacement so she asked if we could connect a spare pump up to the system temporarily until she can send us a couple of spares.  Wired a spare pump up so it can be plugged into the wall and connected it to the system.  The reading on the flow meter for line 3 was reading .4.  Notified the mentor of the situation and will await further instructions and parts.  RTS at 1945 GMT.

8/20/02

MR

The instrument mentor wants the SE25M and SE60M lines checked for blockage and the connectors checked.  Connected a vacuum pump and gauge up to the lines inside the shelter.  I was able to pull a vacuum of 10 psi on both lines and then let it bleed off.  The lines would bleed off in about 2 1/2 to 3 seconds.  I talked with the mentor about this and she agreed that this was about normal.  Pump # 3 was found to be off and had locked up on 8/2/02.  I turned the pump on and it started up and was running fine with the exception of being a little more noisy then the others.  The mentor then had me connect the SE25M and SE60M lines back up in the system to check for airflow.  I will check on the pump in the morning to see if it is still working.  RTS at 2000.

8/5/02

DAN

Performed the work Joe Berry summarizes to Igor Pesenson in an email quoted below.

 

>>> The pump on level 3 froze-up over the weekend.  Dan Nelson has attached the line from level 3 (which has a working filter assembly) to the level 4 pump (which still works). This change occurred at 17:10 on 08/05/02, and the system passed a leak test. Please note that the software now reports level 3 data as level 4. Are you able to change the analysis program or should I change the data logger program to correct the data output? <<<

8/2/02

DAN

Joe asked Site Ops to switch these sample lines (inside the PGS shelter) with the new lines until a repair could be affected on the failed connectors.  When this task was completed, we discovered that the new 60m level was also no longer operational.  Inspection by telescope seems to indicate that the new 60m fixed connector has been bent out of position and is not mating with its other half.  The new 25m sampling point was functioning fine.  This was near 1630 local time, so further repairs were put off until Monday, 8/5/02.

8/1/02

DAN

Significant 60m tower T/Rh probe calibration work was undertaken by David Cook and Mike Ritsche.  Elevator lock out had caused these calibrations to be way behind schedule.  David and Mike’s work required them to lower all elevators to ground level for several hours.  The calibration work was completed and the elevators returned to their operating positions.

 

Some time after raising the elevators, Joe Berry discovered (remotely from LBNL) that we no longer had flow from the SE sample lines at either of the 25m or 60m levels.  When lowered for inspection, we discovered that an o-ring was missing from both connectors.  By process of elimination, we also determined that the missing o-rings must be still in the fixed portion of the connectors, thus not allowing them to fully mate and open their plunger valves to initiate flow.

7/31/02

DAN

Tower Systems, Inc (TSI) came to SGP to inspect and repair the 60m tower damage related to the severe ice storm of late January 2002.  Several fallen u-bolts have been found at the tower base and as a result, elevator operation had been suspended since the ice storm.  During this visit, TSI was also contracted to install a second set of four isotope sampling lines and their associated filters on the west side of the 60m tower for LBNL.  The existing sample lines are located on the SE side of the tower.

 

During preparations for the installation, Site Ops discovered that only enough Dekoron tubing to complete the 60m and 25m levels was available for installation.  Also, TSI observed that the two-part filtration assemblies for these levels were not exact duplicates of those previously installed on the tower.  Following a brief conference, TSI felt that they could make the new, two-piece filter assemblies work.  Site Ops felt that the two lower levels (4m and 2m) could be installed from a ladder when more tubing was available.

 

Work proceeded as planned, with only minor alterations to the fixed portion of the filter assemblies necessary to allow proper clearance from moving tower parts (the elevators).  Flow testing at all phases of the installation appeared normal.  The test results were passed on to LBNL for review.  With all sample line installation (and tower repair) work complete, TSI was released to return home on 7/31/02, late in the afternoon.

7/31/02

CT

Instrument mentor requested that the PGS computer be power cycled. The PGS computer was power cycled at 1357 GMT. The mentor was notified by telephone of the action performed. Status OK. RTS at 1412 GMT.

7/24/02

Mentor

Marc Fisher on site and R/R reference cylinder and adjusted the regulator to 10 psi.

7/23/02

MR

The instrument mentor was notified about the reference cylinder using too much gas. The mentor called back and said there may be a leak and instructed me to turn off the gas on the cylinder.  One of the mentors will be on site on 7/24/02 and will look at the instrument then. Turned off the gas at 1800 GMT.

7/18/02

MR

The instrument mentor requests that the reference flow gas line be checked for leaks.  Checked the line for leaks and none were found.  Checked and tightened all connections.  Adjusted the flow from 38 psi to 15 psi.  Adjusted the needle valve flow from one bubble per second to 3 bubbles per second.  RTS at 1845 GMT.

7/17/02

CT

The instrument mentor requested that the PGS computer be power cycled. The program was hung up. The PGS computer was power cycled at 1707 GMT. Status OK. The mentor was contacted on the status of computer. RTS at 1722 GMT.

7/16/02

CT

The computer program was terminated. The computer was rebooted at 1532 GMT. The mentor was informed of the action taken. RTS at 1547 GMT.

7/16/02

MR

The new reference CO2 tank has arrived and needs to be installed.  R/R the Co2 reference tank.  Notified the mentor.  RTS at 1345 GMT.

7/15/02

CAM

Igor Pesenson notified Site Ops that he lost communications with the computer while remotely downloading data.  T/S found that the computer was locked up and would not respond.  Power cycled the computer at 1811 GMT.  Verified network communications were restored and RTS at 1815 GMT.

7/15/02

CAM

Joe Berry notified Site Ops he was unable to access the PGS computer over the network.  T/S found that the computer was powered on and operational, but PC208 was not running.  Rebooted the computer at 1635 GMT and started PC208 program.  The computer still could not be pinged over the network.  Power cycled the CentreCom hub at 1645 GMT.  Network communications were then restored.  RTS status O.K.

7/12/02

CAM

Instrument mentor notified Site Ops that he was unable to access the PGS computer and that this morning's data collection had failed.  T/S found that the PGS computer could not be pinged or ping out over the network.  All network hardware indicators and connections appeared normal.  Power cycled the hub in the PGS building at 1855 GMT.  Network access resumed successful at 1900 GMT.

7/9/02

CT

The PC208W program was terminated. A Application Error was displayed. "General Protection Fault in module PC208W.EXE."  There was an additional error displayed " WOWEXEC General Protection Fault error Module KRNL 386.EXE." The error application boxes were closed. The PC208W program was restarted at 1540 GMT. RTS at 1555 GMT. The mentor was notified of the corrective maintenance performed.

7/5/02

CT

The PC208W program was terminated.  The message stated "PC208W caused a General Protection Fault in module PC208W.Exe." The computer was rebooted at 1702 GMT. The mentor was notified of the corrective maintenance performed. RTS at 1717 GMT.

7/2/02

CT

The #3 sample flow reading was very high and water was present in the flow meter. The #3 glass bubbling tube in the refrigerator was 3/4 full with water. The mentor was contacted at 1700 GMT.  At the request of the mentor the bubbling tube hose was taken off and cleared of any obstructions. The toggle switch for the #3 sample flow pump was turned to the off position. Waiting on word from the mentor for further corrective maintenance. RTS at 1815 GMT.

7/2/02

CT

The computer needed to be rebooted. The computer was rebooted at 1658 GMT. Status OK. RTS at 1713 GMT.

6/20/02

CAM

Instrument mentor notified Site Ops that the last PGS data available was from 6/18 and that the data did not look good.  Checked the system and found that the PC208 program was open, but was not connected to the data logger.  Connected to the data logger at 1502 GMT and the data began updating.  Checked the IRGA and it did not appear to be on.  There were no lights lit on the front panel.   Switched the front panel power switch off and back on with no change.  Checked the fuse and found it was ok.  Disconnected the power cord for a few seconds and then reconnected.  The LCD backlight and Ready LED lit up briefly and then went off.  The LCD display showed MP in the upper left corner and TEMP in the upper right corner.  Talked with Margaret Torn and she suggested that the power switch needed to be cycled again after the power was disconnected.  Cycled the power switch again at 1750 GMT.  Again, the LCD backlight and Ready LED lit briefly and then went out.  The LCD display now reads This Display Is Not Configured.  It is not clear if the IRGA is working properly.  Also observed that the 16 flask position indicator was in position 1.  Left voice mail for mentor detailing the findings.

6/18/02

MR

The instrument mentor wants the IRGA instrument power cycled. Power cycled the IRGA instrument.  Notified the mentor.  RTS at 1630 GMT.

6/18/02

MR

The instrument mentor wants the refrigerator defrosted.  Shut the air pumps off at 1430 GMT.  Shut off the refrigerator and opened the door to allow the ice to melt.  After 1 hour the ice was just beginning to melt.  Using a heat gun I melted the ice and removed all the water out of the refrigerator.  Turned the refrigerator back on and restarted the air pumps at 1615 GMT.  RTS at 1615 GMT.

6/17/02

CT

The instrument mentor requested that the computer be restarted. The computer was restarted at 1553 GMT. RTS at 1608 GMT.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

6/7/02

CT

The glass tubes in the refrigerator are collecting a build up of condensation. The problem was reported to the mentor by e-mail on 6/6/02. The mentor requested that the refrigerator be turned to its maximum coolest temperature. At 1450 GMT the refrigerator was turned to its maximum coolest temperature. RTS at 1505 GMT.

6/6/02

CT

#1, #2, and #3 glass tubes in the refrigerator are collecting condensation in them. The #4 glass tube has water in it. The mentor has been notified of the situation. Troubleshooting is in progress. RTS at 1500 GMT.

6/6/02

CT

The PC208W program was terminated. The PC208W program was started at 1441 GMT. The mentor has been informed of the problem. RTS at 1456 GMT.

6/6/02

CT

The water in the bubbling tray was overflowing into the refrigerator. The bubbling tray was emptied at 1440. RTS at 1455 GMT.

5/23/02

CT

The PC208W computer program was terminated. The error displayed was " Application Error" The program was closed and then restarted at 1608 GMT. Status OK. RTS at 1623 GMT.

5/16/02

CT

The program was terminated. A application error had occurred. At 1528 GMT the PC208W was closed and then restarted. Status OK. RTS at 1543 GMT.

5/9/02

CT

The computer was found powered off.  At 1633 GMT powered on the computer.  At 1639 GMT after power up, the computer came online and the program began normal operation.  RTS at 1644 GMT.

5/7/02

CT

The bubbling tubes tray that is located inside the refrigerator was overflowing. The bottom of the refrigerator was wet. The refrigerator was wiped dry with towels. The problem was reported to Igor Pesenson by E-mail. RTS at 1601 GMT.

5/7/02

CT

The PC208W program was terminated. The PC208W program was restarted at 1546 GMT. Igor Pesenson was notified by E-mail concerning the problem. RTS at 1601 GMT.

5/2/02

CT

The computer program was terminated. The message displayed was "Application Error"  The program was restarted at 1637 GMT.  RTS at 1652 GMT.  Igor Pesenson was notified of the action taken by E-mail.

4/30/02

CT

The refrigerator inside the PGS building was leaking water on the floor. The Bubbling tubes tray has a possible leak. The tray was removed and replaced. RTS at 1845 GMT.

4/30/02

CT

The computer program had terminated. The Message that was displayed was "Application Error".Igor Pesenson was contacted. The program was restarted at 1830 GMT. RTS at 1845 GMT.

4/25/02

CT

The inside of the refrigerator was wet due to the overflow from the bubbling tubes tray. The level on the Bubbling tubes water tray was drained partially at 1630 GMT. Status OK. RTS at 1645 GMT.

4/25/02

CT

The PC208W program had a application error. At 1635 GMT the PC208W program was closed and restarted. Status OK. RTS at 1650 GMT.

4/24/02

CAM

Instrument mentor notified Site Ops that the PGS computer had locked up on him.  Found that the computer was at the log on screen, but could not log on to the network.  Rebooted the computer and was then able to log on.   RTS at 1820 GMT.

4/22/02

MR

The instrument mentor cannot communicate with the PC208 program and wants the computer rebooted.  Rebooted the computer at 1900 GMT.  RTS at 1915 GMT.

4/18/02

SS

The instrument computer was powered off.  The UPS was found to be bad.  The *UPS was removed and replaced at 1350 GMT.  When the replacement was complete, the computer was powered on and the 208W program was restarted.  The graphical data returned and the instrument computer was operating properly.  RTS 1500 GMT.  Status was OK at this time.   *Old - 5514   *New - 5326

4/17/02

SS

The computer indicated an error message: “208W has caused a stack fault”                                      No graphical data was being plotted.  Restarted 208W at 1335 GMT.                                                         When the restart was complete, the data began to plot and the program                                                        appeared to be running properly.  RTS 1350 GMT.  Status was OK at this                                                  time.

4/16/02

SS

An error message occurred on the instrument computer.  Restarted 208W                                                  at 1355 GMT.  When the restart was complete, the program began to run                                                   properly and the graphical data returned.  RTS 1410 GMT.  Status was                                                      OK at this time.

4/15/02

CT

Instrument mentor requested that the PGS computer be rebooted. At 1345 GMT the computer was power cycled. Status OK. RTS at 1400 GMT.

4/11/02

SS

An error message was present on the computer.    PC208W was closed and restarted at 1335 GMT.  When the restart was complete the program began to run properly.  RTS 1350 GMT.  Status was OK at this time.

4/9/02

SS

The mentor indicated that the instrument computer needed to be powered on.  The computer was powered on at 1715 GMT.  The computer and program then began to operate properly.  RTS 1730 GMT.  Status was OK at this time.

3/28/02

CAM

Instrument mentor requested that the reference cylinder low pressure be increased to 12 psi.  Adjusted the pressure from 9 psi to 12 psi at 2004 GMT.

3/27/02

SS

Instrument mentor reported a problem with the reference gas.  Found that the low side pressure was set to near zero.  Mentor suggested setting the pressure be set to what the other tanks are reading.  Set the pressure to 10 psi at 2210 GMT.

3/26/02

SS

The reference cylinder pressure was reading 0psi.  Removed and replaced the reference cylinder at 1910 GMT.  When the replacement was complete, the new cylinder pressure was reading 2090psi.  Status was OK at this time.

3/21/02

CAM

Check of PGS system following a hard power bump found that the computer was powered off.  Powered the computer on at 1946 GMT.  Started PC208 software and RTS status O.K. at 1947 GMT.

3/20/02

CAM

Sample line 3 was not blowing bubbles in the water drain pan.  Checked pump #3 and it was operational.  Disconnected the stainless steel line from the bottom of the glass tube on sample line 3 and observed that air was blowing out of the tube.  There was a small piece of what appeared to be plastic stuck in the stainless steel line opening.  Removed the debris and reconnected the line to the tube.  Sample line 3 is now bubbling in the drain pan as it should.  RTS at 1630 GMT.

3/15/02

KT

Mentor stated that he could not FTP into the CO2 FLUX computer.  At 1932 GMT observed that the UPS that the CO2 FLUX computer was plugged into had no power and found that the computer was off.  At 1937 GMT after power cycling the UPS, the computer came online and gave an ERROR 2 message.  The graph was not scrolling and the UPS was giving an A8 alarm.  After conversing with mentor, the CO2 FLUX computer was disconnected from the UPS and was plugged into the power strip to be left as is over the weekend.  The CO2 FLUX computer was left online and mentor was able to FTP to system.  RTS at 1955 GMT.

3/12/02

CAM

Instrument mentor advised that the IRGA power did not appear to be restored following the weekend outage.   The unit requires a manual reset after a long power outage.  Opened the box in the electronics rack and reset the power switch on the IRGA.  RTS status O.K. at 2058 GMT.

3/11/02

KT

The UPS for the 60m flux computer needed to be replaced.  At 1933 GMT removed and replaced the *UPS.  At 1940 GMT after installation, the computer was powered on and the program began running and the graph was updating.  RTS at 1945 GMT.   *Old - LI360V05323  *New - LI660VB21645

3/11/02

MR

The power to the site was lost over the weekend.  The instrument was found to be off.  Powered up the computer at 1450 GMT.  System came up and started taking data at 1454 GMT.  RTS at 1505. 

3/7/02

CT

Instrument mentor requested trouble shooting on the bubble tubes in the drying sample refrigerator. At 1843 GMT mentor requested that suction be applied to check out any blockage. Troubleshooting came up negative for blockage. RTS at 1854 GMT.

3/4/02

CAM

Instrument mentor notified Site Ops that he was unable to connect to the PGS computer.  Checked on the system and found that the computer was powered off.  Powered the computer on and RTS at 1755 GMT.  The site power was out for 1.5 hours on 3/2 and was responsible for the shut down.

2/13/02

MR

The instrument mentor needs the CO2 Sampling system controller power cycled due to it has never recovered since the power outage on 1/30/02.  Power cycled the system at 1455 GMT.  RTS at 1510 GMT.

2/12/02

MR

The AC power to the site was lost on 1/30/02 due to local ice storm.  Powered up the system at 1508 GMT.  The system seemed to power up ok.  Notified Margaret Torn.  Status ok.  RTS at 1523 GMT.

1/29/02

CT

The reference cylinder has exceeded the expected usage amount. The problem has been reported to the mentor. RTS at 1705 GMT.

1/28/02

TG

Marc Fisher requests that FTP Server be installed on the PGS computer.  Technician stopped the PGS software at 2005 GMT. Technician installed IE 5.5 and NT 4.0 Option Pack (FTP Server). System restarted at 2121 GMT.

1/24/02

MR

The mentor notified me that the flasks were ready to be removed and shipped.  Closed the stopcocks and removed the flasks.  Installed the next set into the cooler and readied them for sampling.  Notified the mentor.

1/22/02

MR

The mentor notified us that the CO2 Flasks need to be loaded and ready so they can run samples this week.  The flasks were already loaded into the cooler.  Checked the numbers on the paper with the flask numbers and opened all stopcocks.  Notified the mentor.

1/22/02

CT

The #3 graph is not graphing. The mentor has been notified by email of the current graphing problem. No further corrective action has been taken at this time. RTS at 1635 GMT.

1/10/02

CJ

The computer screen display was blank. Mike Rainwater rebooted computer at 1954 GMT. Status OK. RTS at 2009 GMT.

12/20/2001

CJ

Graph #3 does not appear to be plotting the data. 12/20/2001 mentor has been notified by email of the problem.

12/19/01

MR

The tower needs to be lowered due to calibration on the PGS LICOR system. Lowered the tower at 2055 GMT.  Performed the calibration.  Raised the tower at 2110 GMT.

12/18/2001

CJ

ST420 cylinder pressure decrease is out of the expected consumption range. Problem was reported to mentor.

12/18/2001

CJ

ST360 cylinder pressure decrease is out of the expected consumption range. Problem was reported to mentor.

12/13/2001

CJ

PC208W did not appear to be running. Restarted PC208W at 1608 GMT. RTS at 1623 GMT.

12/06/01

KT

The PC208w program was found to not be running.  At 1710 GMT started PC208w.  At 1712 GMT the program came on line and after checking the graphs, status ok.  RTS at 1723 GMT.

11/29/01

CJ

ST420 has used more than 25 PSI since last reading. No further CM at this time.

11/27/01

CAM

The PC208 program was running but was not connected to the datalogger and no graphs were displayed.  Connected to the data logger at 1809 GMT and the graphs were then displayed.  Graphs 1 and 2 are displaying reasonable data, but graph 3 displays all zeros. 

11/26/01

KT

The mentor requested that the REF gas tank be changed out.    At 2025 GMT the empty tank was removed and the new full tank was installed.  After installation the gauge readings were,  2000 PSI for the tank pressure,  and 10 for the REF reading.  RTS at 2105 GMT. 

11/26/01

CAM

The PC208 program was running, but was not connected to the data logger.  Clicked on connect at 1849 GMT.  The graphical displays 1-3 opened and 1-2 began updating.

11/20/01

CAM

Heather Cooley notified Site Ops that the PGS computer could not be accessed over the network at 1945 GMT.  Check of system found the PGS computer was powered off.  The site power was shut off from 1503 to 1555 GMT today for trench work and the PGS had not restarted.  Powered the computer on and started PC208W at 2032 GMT.  The program started normally, but all data values were reporting 0’s on the display.  Passed this information along to the mentor.

11/05/2001

TMG

Technician removed D4 program from startup menu. Technician rebooted computer and RTS at 2127 GMT.

11/5/01

CAM

Sample lines 1 and 2 were showing no air flow.  T/S found that pumps 1 and 2 have failed.  The pumps were extremely hot.  The exterior cooling fan that was attached to the box has also failed.  Powered the pumps off at 2100 GMT.   The instrument mentor will be on site next week and will bring replacement pumps.  Removed the failed fan and will replace with a spare.  System RTS with only pumps 3 and 4 running. 

11/01/2001

CJ

At 1730 GMT graphs are showing zero values for data. Mentor will be notified. RTS at 1745 GMT.

10/30/01

CJ

Program found locked up at 1645 GMT.  Restarted program at 1645 GMT. RTS at 1700 GMT.

10/25/01

CM/TG

The PGS time was 9 minutes and 21 seconds ahead of SDS time.  Installed D4 program on the PGS computer and configured it to sync the clock to SDS every 15 minutes.  Updated the time at 1845 GMT.  Set the data logger clock to PC time at 1847 GMT.  RTS status O.K. at 1850 GMT.

10/4/01

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1800 GMT.  Raised the tower at 1820 GMT.  RTS at 1830.

10/3/01

MR

The tower was lowered due to calibration on other equipment causing data interruption.  Lowered the tower at 2105 GMT.  Raised the tower at 2155.  RTS at 2205.

10/2/01

MR

The tower had to be lowered due to other maintenance.  Lowered the tower at 1548.  Raised the tower at 1600.  RTS at 1603.

9/27/01

MR

Lowered the 25 and 60 meter tower due to maintenance being performed on the 25 meter tower.  Lowered the tower at 1503.  Raised the tower at 1912.

8/29/01

MR

The instrument mentor has sent a procedure to check and reset the valve controller. Checked the valve controller and it would not advance past the #10 position.  Reset the controller as per Larry Giles instruction by depressing he STEP and HOME keys at the same time and then pressing the STEP switch to switch the controller to the 16 position.  Cycled through the 16 valve positions and then set the position to #8 and pressed the HOME key.  The unit returned to position #1 and is ready to take samples.  Notified Margaret Torn and she said she would set the instrument up to take samples starting on Wednesday night and wants the instrument checked on Thursday.  RTS at 1800 GMT.

8/20/01

CAM

The site experienced a 3.5 hour power outage over the weekend, beginning at 0033 GMT on 8/18.  As a result, the PGS computer shut down and did not restart.   Powered the PGS computer on at 1740 GMT on 8/20.  RTS status O.K.

8/14/01

JEM

At 1500 GMT while checking instruments and data noticed that the PC10x datalogger computer display was not graphing collected data. At 1900 GMT performed a reboot of the computer. The computer display was showing graphed data and was operational at 1915 GMT.

8/7/01-8/8/01

MR

Southeast booms lowered at 1918 GMT on 8/7 for 25M and 60M T/H calibrations.  Booms were left down at ground level overnight.  Raised booms and RTS at 1840 GMT on 8/8.

7/20/01

MR

Lowered the tower at 1450.  Performed a Licor calibration procedure.  Raised the tower at 1700 GMT.

7/5/01

CAM

During the weekly PM checks the technician found that the computer was displaying a system kernel error.  Technician OK'd the error and rebooted the computer.  The computer came back up with the same kernel error.  Checked on the system at 2030 GMT and found that the error was cleared and the computer appeared normal but PC208W was not running.  Started PC208W program at 2045 GMT and RTS status O.K.

6/29/01

CAM

Instrument mentor notified Site Ops that she could not access the PGS computer over the network.  Checked the PGS status and found that the program was running and appeared to be OK.  Could also successfully ping the computer over the network from a local machine.  Restarted Timbuktu program at 2030 GMT just to be sure there wasn’t a problem with it.  Notified mentor of the findings and she later reported she was able to access the computer.  Larry Giles had been on site earlier in the day performing maintenance on the PGS system.  Suspect the computer may have been off line when the mentor initially tried to access it.

6/18/01

KT

Mentor requests that a *UPS be installed and the power strip that has the PGS computer, the Fiber Optic Transceiver and the network HUB be plugged into the UPS.  At 2039 GMT powered off the PGS computer.  At 2041 GMT plugged UPS into wall outlet.  At 2044 GMT plugged the PGS computer, the Fiber Optic Transceiver and the network HUB into the back the UPS.  At 2045 GMT powered UPS on.  At 2048 GMT powered the PGS computer on.  RTS at 2054 GMT.  *New – LI360V05514

6/18/01

CAM

Instrument mentor notified Site Ops that the PGS computer was not accessible over the network. T/S found that the computer and hub were without power. The UPS that was installed last week had failed over the weekend. The site experienced several power bumps and the UPS was unable to keep the computer running due to weak batteries. Removed the *UPS and plugged the computer and hub to straight AC power. Powered the computer on and restarted PC208W at 1635 GMT. Will install another UPS when another spare becomes available or the batteries are replaced in the old one. *Old - LI360V05712

6/15/01

CAM

The PGS computer was not connected to UPS power. This caused the computer to shut down whenever a power bump occurred. Located a spare *UPS and installed it in the building. Shut the PGS computer down at 1657 GMT. Transferred the computer, hub, and FOT power cables to the UPS and powered everything back up at 1700 GMT. Restarted PC208W and RTS at 1702 GMT. *New - LI360V05712

6/14/01

CAM

Instrument mentor notified Site Ops that the PGS computer was not online. Neither the PGS nor the CO2 Flux computer could be accessed over the network. Status check found that the PGS computer was powered off. Powered the computer on at 1429 GMT and started the program. Attempts to ping either computer were still unsuccessful. Power cycled the hub and the FOT at 1435 GMT and the computers were then accessible over the network. RTS status O.K. at 1444 GMT. The problem is that the PGS computer and hub are not connected to UPS power. This causes the computer to shut down whenever the building experiences a power bump. Will locate a spare UPS and install it in the building to supply power to the PGS computer and hub during power bumps.

6/1/01

CAM

Instrument mentor notified Site Ops that the PGS computer was not online. Instrument status check found that the computer was powered off. The site experienced an extended power outage during thunderstorms on 5/30. Powered the computer on at 1309 GMT. Checked the BIOS settings to see if the PWRON After PWR-FAIL option was enabled. The option was enabled but the computer still will not auto-start after a power outage. Rebooted the computer and RTS at 1319 GMT. Checked PC208W directory for a raw data file but there was none. The last file transferred was PGS052801. RTS at 1320 GMT.

5/18/01

TG

Instrument mentor could not perform remote control on computer. T/S of system indicates that the 10/100 hub is bad. R/R hub and system RTS status O.K. at 2140 GMT.

5/18/01

CAM

The pump for sample line #1 has failed. Mentor is aware of this and will send a spare. Powered the pump off and closed the valve in the refrigerator at 2130 GMT.

5/18/01

MR

The computer was found off. Turned on the computer. RTS at 1610. Line 1 is not bubbling and sample one is reading 0 flow. Called the mentor and left a message.

5/8/01

CAM

The PC clock was set to CST and was one hour behind due to daylight savings. The clock needed to be set to GMT. Closed all programs and logged on as administrator at 2028 GMT. Set the PC time to GMT at 2035 GMT. Shut down and logged on as operator at 2038 GMT. Restarted PC208W and RTS at 2040 GMT. Notified mentor that all data values on the graphs are zero. Awaiting reply.

5/8/01

CAM

Margaret Torn requested that the PGS computer be rebooted and PC208W started. Rebooted the computer at 1708 GMT and started the PC208W program. RTS at 1710 GMT. The PC clock is set to CST and is one hour behind due to daylight savings. Mentor requested that the clock be set to GMT, however this requires to logon with administrative privileges. The system is downloading data right now, so will change the time later.

5/4/01

MR

Performed the weekly. The gauges read as follows; ST330=1825, ST360=1825, ST390=1749, ST420=1825, REF=1375. Status checks O.K.

5/01/01

MR

Performed the weekly. The gas cylinders readings where as follows; ST330=1825, ST360=1825, ST390=1750, ST420=1825, Ref=1400. Status O.K.

4/27/01

MR

Performed the first weekly for the system. The air gauge readings were as follows: ST330=1825, ST360=1825, ST390=1750, ST420=1825, REF=1450. Line 2 in the refrigerator had water in the tube. Removed the water and readjusted the air flow through the tube so that it was bubbling the same as the other 3. All other readings and values for the weekly where normal. No further PM at this time.

4/23/01

KT

Instrument mentor requested that the #3 line be disconnected and connected back up to the testing line. At 2000 GMT disconnected the #3 line from the back of the enclosure and connected it up to the testing line as was instructed by the technicians who were on site when instrument was installed. At 2007 GMT upon change out, the #3 cylinder valve was turned on. System line back in testing phase until further notice. RTS at 2015 GMT.

4/20/01

MR

Performed the purge gas flow test. Started the test at 1700 GMT. Stopped the test at 2000. Unhooked line #3 from the air tank and connected it back up to the instrument as instructed by the mentors last week. Turned tank #3 off and placed a cap on the line.

4/19/01

CAM

Received confirmation from mentor that the drain on line 2 should be open in the refrigerator. The valve did not appear to be completely closed, but was not open enough to allow the water to be blown out. Opened the valve at 1645 GMT enough so that all the water drained out of the glass tube into the pan. Then closed the valve just to the point that it was blowing the same amount of air as the others.

4/19/01

CAM

Tim Grove requested that the PC be rebooted due to mentor experiencing problems running Timbuktu yesterday. Rebooted the computer at 1355 GMT and RTS at 1400 GMT. Opened the refrigerator door for a visual inspection and found that the glass tube on line 2 was full of water. The drain valve on that line was closed, therefore not allowing the water to escape out the bottom. Sent email to mentor requesting confirmation that the valve should be open during this first week of operation.

4/19/01

CAM

Mentor requested that the valve on the tank marked Air 8 be opened for 3 hours each day this week. Opened the valve on the tank at 1352 GMT and closed at 1652 GMT.

4/18/01

CAM

Mentor requested that the valve on the tank marked Air 8 be opened for 3 hours each day this week. Opened the valve on the tank at 1338 GMT and closed at 1647 GMT. Final tank 8 pressure reading 1025. Closed the tank main valve and notified mentor of the large difference in pressure readings between yesterday and today.

4/17/01

CAM

Mentor requested that the valve on the tank marked Air 8 be opened for 3 hours each day this week. Opened the valve on the tank at 1345 GMT and closed at 1643 GMT. Final tank 8 pressure reading 1720.

4/16/01

CAM

Mentor requested that the valve on the tank marked Air 8 be opened for 3 hours each day this week. Opened the valve on the tank at 1350 GMT and closed at 1651 GMT. Final tank 8 pressure reading 1725.

4/9/01 -4/13/01

Mentor

Precision Gas Sampling (PGS) system installed at the 60 M tower. The instruments and computer are located in a portable building at the tower base.